### v1

# Asymptotic estimates for solutions of inhomogeneous non-divergence diffusion equations with drifts

Abstract.
We study the long-time dynamics of the nonlinear processes modeled by diffusion-transport partial differential equations in non-divergence form with drifts. The solutions are subject to some inhomogeneous Dirichlet boundary condition. Starting with the reduced linear problem, we obtain the asymptotic estimates for the solutions, as time $t\to\infty$, depending on the asymptotic behavior of the forcing term and boundary data. These are established in both cases when the drifts are uniformly bounded, and unbounded as $t\to\infty$. For the nonlinear problem, we prove the convergence of the solutions under suitable conditions that balance the growth of the nonlinear term with the decay of the data. To take advantage of the diffusion in the non-divergence form, we prove an inhomogeneous version of the Landis-typed Growth Lemma and apply it to successive time-intervals. At each time step, the center for the barrier function is selected carefully to optimize the contracting factor. Our rigorous results show the robustness of the model.
Key words and phrases:
diffusion-transport, fluids in porous media, nonlinear PDE, non-divergence equation, qualitative study, Bernstein–Cole–Hopf, asymptotic analysis
2020 Mathematics Subject Classification:
35Q35, 35A09, 35B50, 35B40
Corresponding author.
1. Introduction
We consider a fluid flow in porous media in the $n$-dimensional Euclidean space. (For physical problems, $n\leq 3$). Let $x\in\mathbb{R}^{n}$ be the spatial variable and $t\in\mathbb{R}$ the time variable. Let $\rho(x,t)$, $p(x,t)$ and $v(x,t)$ denote the density, pressure and velocity of the fluid at location $x$ and time $t$. In [8], we derived an equation based on Eistein’s probabilistic method for the Brownian motion [6], namely,
$\frac{\partial\rho}{\partial t}=\langle A(x,t),D^{2}\rho\rangle+(M_{0}(x,t)v(x,t))\cdot\nabla\rho,$ (1.1)
where $A(x,t)$ and $M_{0}(x,t)$ are $n\times n$ matrices such that
$\xi^{\rm T}A(x,t)\xi>0\text{ and }\xi^{\rm T}M_{0}(x,t)\xi\geq 0\text{ for all}\xi\in\mathbb{R}^{n}.$
In equation (1.1), the term $\langle A(x,t),D^{2}\rho\rangle$ denotes the inner product between the matrix $A(x,t)$ and the Hessian matrix $D^{2}\rho$, see the notation in section 3 below. It represents the diffusion in the non-divergence form, while the term $(M_{0}(x,t)v(x,t))\cdot\nabla\rho$ represents the transport.
Assume the anisotropic Darcy’s law [5, 1],
$v=-\bar{K}(x,t)(\nabla p-\rho\vec{g}),$ (1.2)
where $\bar{K}(x,t)$ is an $n\times n$ matrix, and $\vec{g}$ is the gravitational acceleration in the case $n=1,2,3$ or any constant vector otherwise. Combining (1.1) with (1.2) yields
$\frac{\partial\rho}{\partial t}=\langle A(x,t),D^{2}\rho\rangle-(K_{0}(x,t)\nabla p)\cdot\nabla\rho+\rho B_{0}(x,t)\cdot\nabla\rho,$ (1.3)
where
$K_{0}(x,t)=M_{0}(x,t)\bar{K}(x,t),\quad B_{0}(x,t)=M_{0}(x,t)\bar{K}(x,t)\vec{g}.$
When additional relations between $\rho$ and $p$ are specified, we can derive from (1.3) a nonlinear partial differential equation (PDE) for $\rho$. The following are some typical cases in fluid mechanics.
Case of isentropic gas flows. In this case, $p=c\rho^{\gamma}$ for some constant $c>0$, where $\gamma\geq 1$ is the specific heat ratio. Then equation (1.3) has the specific form
$\frac{\partial\rho}{\partial t}=\langle A(x,t),D^{2}\rho\rangle-c\gamma\rho^{\gamma-1}(K_{0}(x,t)\nabla\rho)\cdot\nabla\rho+\rho B_{0}(x,t)\cdot\nabla\rho.$ (1.4)
In particularly, one has $\gamma=1$ for ideal gases and equation (1.4) reads as
$\frac{\partial\rho}{\partial t}=\langle A(x,t),D^{2}\rho\rangle-c(K_{0}(x,t)\nabla\rho)\cdot\nabla\rho+\rho B_{0}(x,t)\cdot\nabla\rho.$ (1.5)
Case of slightly compressible fluids. The equation of state is
[…]
Then (1.3) can be rewritten as
$\frac{\partial\rho}{\partial t}=\langle A(x,t),D^{2}\rho\rangle-\frac{1}{\kappa\rho}(K_{0}(x,t)\nabla\rho)\cdot\nabla\rho+\rho B_{0}(x,t)\cdot\nabla\rho.$ (1.6)
Equations (1.4), (1.5) and (1.6) fall into a larger class of equations which is formulated by (2.3). This nonlinear parabolic equation in non-divergence form will be the subject of our study. To the best of our knowledge, there are not many papers dealing with the long-time dynamics of non-divergence parabolic equations, especially the highly nonlinear ones such as (2.3). (However, see [14] for results for non-divergence linear elliptic problems with unbounded drift.) In our previous work [8], we essentially study (2.3) with zero forcing and, after shifting the solution by a constant, the homogeneous boundary condition. The purpose of the current paper is to obtain results for inhomogeneous problems. We emphasize that the equations we are dealing with contain drifts which can be nonlinear and unbounded as time tends to infinity. As we will see, it requires more technical and sophisticated tools.
There are three ingredients that we develop in the current paper to study of this highly nonlinear problem. The first is the general Bernstein–Cole–Hopf transformations that convert the equation to more manageable inequalities. The second is the Growth Lemma for inhomogeneous problems expressed in terms of the time-integral of the forcing function. The third is the iteration of this lemma on successive time-intervals with appropriate adjustments at each step. With those, we are able to establish the convergence/decay of the solutions as time tends to infinity.
The paper is organized as follow. In section 2, we recall the notation, set the background and introduce the main initial boundary value problem (2.9). In section 3, we use the standard Maximum Principle to derive preliminary bounds for the solutions and sub-/super-solutions of inhomogeneous problems for both linear and nonlinear equations. A new feature compared to the usual form in literature is that the bounds are expressed in terms of the integral in time for the forcing function. This hints at estimates that we will obtain and explore later which are similar to the ones for the two-dimensional Navier–Stokes equations by Foias and Prodi [7]. In section 4, we study the linear problem. We recall in Lemma 4.3 the Growth Lemma in a finite interval for the homogeneous problem. This is extended to the inhomogeneous problem in Lemma 4.4. A general but complicated estimate for all time is established in Proposition 4.5. The main idea is combining the application of the Growth Lemma in successive intervals of time with the moving center in the construction of the barrier function at each step. This allows us to gain better contracting factors of the Growth Lemma (not too close to $1$.) The estimate in Proposition 4.5 is then used to establish a particular decay estimate for the solution in the case of bounded drifts, see Theorem 4.7. The case of unbounded drifts as $t\to\infty$ is treated in Theorems 4.11 and 4.14. Basically, even if the drift term goes to infinity as $t\to\infty$, we still have the boundedness or even the decay for the solution if this growth is not too fast and the forcing function and boundary data decay at an appropriate rate to balance out that growth. Although the assumptions are technical, some explicit and simple examples are provided in Propositions 4.15 and 4.16 to demonstrate the applicability of the results. In section 5, we study the nonlinear problem. We recall from our previous work [8] the transformation of Bernstein–Cole-Hopf type [2, 4, 11] in order to remove the quadratic term of the gradient. However, unlike [10] the resulting partial differential inequalities (5.4) and (5.6) are still highly nonlinear. Nonetheless, they will be treated by the techniques developed in the preceding section for linear equations with possible unbounded drifts at time infinity. The case for $L^{1}_{t}$-forcing function is treated in Theorems 5.6 and 5.7. Note that even when the solution is proved to be bounded, the contributing drift term still may not be so, see (5.38). For the case when the forcing function is not in class $L_{t}^{1}$, we are able to treat the slightly compressible fluids in Theorems 5.9 and 5.11.
It is worth mentioning that the obtained results show the robustness of the dynamical system generated by the studied PDE. This in part justifies our models (1.1)–(1.6) of the filtration for a class of fluid flows in porous media.
2. Preliminaries
Throughout the paper, the spatial dimension $n\geq 1$ is fixed. For a vector $x\in\mathbb{R}^{n}$, its Euclidean norm is denoted by $|x|$. Let $\mathcal{M}^{n\times n}$ denote the set of $n\times n$ matrices of real numbers, and $\mathcal{M}^{n\times n}_{{\rm sym}}$ denote the set of symmetric matrices in $\mathcal{M}^{n\times n}$. For two matrices $A,B\in\mathcal{M}^{n\times n}$, their inner product $\langle A,B\rangle$ is the trace ${\rm Tr}(A^{\rm T}B)$. For a real-valued function $f(x)$ with $x=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}$, we denote by $D^{2}f$ the Hessian matrix of the second partial derivatives $(\partial^{2}f/\partial x_{i}\partial x_{j})_{i,j=1,\ldots,n}$.
For a real-valued function $f$, we recall that the positive and negative parts of $f$ are $f^{+}=\max\{0,f\}$ and $f^{-}=\max\{0,-f\}=(-f)^{+}$. Then one has
$f=f^{+}-f^{-},\ |f|=f^{+}+f^{-}=\max\{f^{+},f^{-}\},$
$0\leq f^{+}\leq|f|,\ 0\leq f^{-}\leq|f|,\text{ and }-f^{-}\leq f\leq f^{+}.$
If a function $f:S\to\mathbb{R}$ is bounded from above, then $\max\Big{\{}0,\sup_{S}f\Big{\}}=\sup_{S}(f^{+})$.
For the remaining of the paper, $U$ is a non-empty, open, bounded subset of $\mathbb{R}^{n}$ with boundary $\Gamma$ and closure $\bar{U}$. Denote $d={\rm diam}(U)=\max\{|x-x^{\prime}|:x,x^{\prime}\in\bar{U}\}.$ For any point $y\not\in\bar{U}$, set
$r_{*}(y)=\min\{|x-y|:x\in\bar{U}\}\text{ and }R_{*}(y)=\max\{|x-y|:x\in\bar{U}\}.$
Then one has, for any $y\not\in\bar{U}$,
$0<R_{*}(y)-d\leq r_{*}(y)<R_{*}(y)\leq r_{*}(y)+d.$ (2.1)
We denote by $C_{x,t}^{2,1}(U\times I)$, for any interval $I$ of $\mathbb{R}$, the class of continuous functions $u(x,t)$ on $U\times I$ with continuous partial derivatives $\partial u/\partial t$, $\partial u/\partial x_{i}$, $\partial^{2}u/\partial x_{i}\partial x_{j}$ for $i,j=1.\ldots,n$.
For a number $T\in(0,\infty)$, denote $U_{T}=U\times(0,T]$ and the parabolic boundary $\Gamma_{T}=\overline{U_{T}}\setminus U_{T}$, where $\overline{U_{T}}=\bar{U}\times[0,T]$ is the closure of $U_{T}$ (in $\mathbb{R}^{n+1}$). For the sake of convenience, we also denote $U_{\infty}=U\times(0,\infty)$ and its boundary $\Gamma_{\infty}=\partial(U_{\infty})$.
Hereafter, $J$ is an interval in $\mathbb{R}$ with non-empty interior, and $P$ is a function in $C^{1}(J,\mathbb{R})$ with the derivative
$P^{\prime}\in C(J,[0,\infty)).$ (2.2)
For fluid flows in porous media, $P$ is related to the equation of state for $\rho$ and $p$. However, we will consider a general function $P$. It is clear from (2.2) that $P$ is an increasing function on $J$.
Let an extended number $T\in(0,\infty]$ be fixed. Let $A:U_{T}\to\mathcal{M}^{n\times n}_{{\rm sym}}$, $K:U_{T}\to\mathcal{M}^{n\times n}$ and $B:U_{T}\times J\to\mathbb{R}^{n}$ be given functions.
We study the nonlinear equations (1.4), (1.5) and (1.6) in the general form
$\frac{\partial u}{\partial t}-\langle A(x,t),D^{2}u\rangle+B(x,t,u)\cdot\nabla u+P^{\prime}(u)(K(x,t)\nabla u)\cdot\nabla u=f(x,t),$ (2.3)
where $f:U_{T}\to\mathbb{R}$ is a given function.
For example, for slightly compressible fluids and equation (1.6), we can use
$J=(0,\infty),\quad P(s)=\ln s,\quad K(x,t)=\kappa^{-1}K_{0}(x,t),\quad B(x,t,s)=-sB_{0}(x,t).$ (2.4)
For isentropic, non-ideal gas flows and equation (1.4) with $\gamma>1$, we can use
$J=[0,\infty),\quad P(s)=s^{\gamma},\quad K(x,t)=cK_{0}(x,t),\quad B(x,t,s)=-sB_{0}(x,t).$ (2.5)
For ideal gas flows and equation (1.5), we can still use (2.5) with $\gamma=1$ for physical $u=\rho$ (the density), but can also alter $J=\mathbb{R}$ in (2.5) for mathematical $u$. Thus,
$J=[0,\infty)\text{ or }J=\mathbb{R},\quad P(s)=s,\quad K(x,t)=cK_{0}(x,t),\quad B(x,t,s)=-sB_{0}(x,t).$ (2.6)
Remark 2.1.
Although we focus on equation (2.3) with the constraint (2.2) in this paper, our obtained results below will apply to a more general class of equations, namely,
$\frac{\partial u}{\partial t}-\langle A(x,t),D^{2}u\rangle+B(x,t,u)\cdot\nabla u+\mu(u)(K(x,t)\nabla u)\cdot\nabla u=f(x,t),$ (2.7)
where $\mu$ is a continuous function from $J$ to $[0,\infty)$ or to $(-\infty,0]$. Indeed, in the former case, we just take $P$ such that $P^{\prime}=\mu$, while in the latter case, we take $P^{\prime}=-\mu$ and the new $K:=-K$. Then equation (2.7) is of the form (2.3) with the function $P$ satisfying (2.2).
We define the nonlinear operator $L$ associated with the left-hand side of (2.3) explicitly as
$Lu=\frac{\partial u}{\partial t}-\langle A(x,t),D^{2}u\rangle+B(x,t,u)\cdot\nabla u+P^{\prime}(u)(K(x,t)\nabla u)\cdot\nabla u$ (2.8)
for any function $u\in C_{x,t}^{2,1}(U_{T})$ with $u(U_{T})\subset J$. The main task of this paper is to study the following initial boundary value problem
$\begin{cases}Lu(x,t)=f(x,t)&\text{ in }U\times(0,\infty),\\ u(x,t)=g(x,t)&\text{ on }\Gamma\times(0,\infty),\\ u(x,0)=u_{0}(x)&\text{ in }U,\end{cases}$ (2.9)
where the forcing term $f:U_{\infty}\to\mathbb{R}$, the boundary data $g:\Gamma\times(0,\infty)\to\mathbb{R}$ and the initial condition $u_{0}:U\to\mathbb{R}$ are given functions.
A function $u(x,t)$ is said to be a solution of Problem (2.9) if it belongs to $C_{x,t}^{2,1}(U_{\infty})$ with $u(U_{\infty})\subset J$ and satisfies all equations in (2.9).
In order to study (2.9), we need some fine properties of solutions or sub-/super-solutions of a related linear inhomogeneous, non-divergence equations with general drifts. It involves the general linear operator of the following form. Given a function $b:U_{T}\to\mathbb{R}^{n}$, we define the linear operator $\widetilde{L}$ by
$\widetilde{L}w=w_{t}-\langle A(x,t),D^{2}w\rangle+b(x,t)\cdot\nabla w\text{ for }w\in C_{x,t}^{2,1}(U_{T}).$ (2.10)
3. Direct estimates by the Maximum Principle
Below, we establish a maximum principle for solutions of the inhomogeneous problem $\mathscr{L}u(x,t)=f(x,t)$ with either nonlinear operator $\mathscr{L}=L$ in (2.8) or the linear one $\mathscr{L}=\widetilde{L}$ in (2.10).
Theorem 3.1.
Given a number $T\in(0,\infty)$, assume that there exists a constant $c_{0}>0$ such that
$\xi^{\rm T}A(x,t)\xi\geq c_{0}|\xi|^{2}\text{ for all $(x,t)\in U_{T}$ and all$\xi\in\mathbb{R}^{n}$.}$ (3.1)
Let $u\in C(\overline{U_{T}})\cap C_{x,t}^{2,1}(U_{T})$ and define, for $t\in(0,T]$,
$G_{1}(t)=\max_{\Gamma_{t}}u,\quad G_{2}(t)=\min_{\Gamma_{t}}u,\quad G(t)=\max_{\Gamma_{t}}|u|.$
Consider the operator $\mathscr{L}$ to be either
  1. ()
    $\mathscr{L}=\widetilde{L}$, or
  2. ()
    $\mathscr{L}=L$ with $u(U_{T})\subset J$.
Then the following statements hold true.
  1. ()
    Suppose there is a functions $F_{1}\in C([0,T],[0,\infty))$ such that $\mathscr{L}u(x,t)\leq F_{1}(t)$ for all $(x,t)\in U_{T}$. Then
    $\max_{\overline{U_{t}}}u\leq G_{1}(t)+\int_{0}^{t}F_{1}(\tau){\rm d}\tau\text{for all }t\in(0,T].$ (3.2)
  2. ()
    Suppose there is a function $F_{2}\in C([0,T],[0,\infty))$ such that $\mathscr{L}u(x,t)\geq-F_{2}(t)$ for all $(x,t)\in U_{T}$. Then
    $\min_{\overline{U_{t}}}u\geq G_{2}(t)-\int_{0}^{t}F_{2}(\tau){\rm d}\tau\text{for all }t\in(0,T].$ (3.3)
  3. ()
    Suppose there is a function $F\in C([0,T],[0,\infty))$ such that $|\mathscr{L}u(x,t)|\leq F(t)$ for all $(x,t)\in U_{T}$. Then
    $\max_{\overline{U_{t}}}|u|\leq G(t)+\int_{0}^{t}F(\tau){\rm d}\tau\text{ for all }t\in(0,T].$ (3.4)
Proof.
We consider Case ()3.1 first.
()3.1 Define, for $t\in[0,T]$, $\mathcal{F}_{1}(t)=\int_{0}^{t}F_{1}(t){\rm d}\tau.$ Set $v(x,t)=u(x,t)-\mathcal{F}_{1}(t)$. Then
$\widetilde{L}v(x,t)=\widetilde{L}u(x,t)-\mathcal{F}_{1}^{\prime}(t)=Lu(x,t)-F_{1}(t)\leq 0.$
On $\Gamma_{t}$, one has $v(x,t)\leq u(x,t)\leq G_{1}(t)$. By the Maximum Principle for $\widetilde{L}$ and $v$, we have $v(x,t)\leq G_{1}(t)$ on $U_{T}$. Thus, $u(x,t)\leq G_{1}(t)+\mathcal{F}_{1}(t)$ on $U_{T}$, which proves (3.2).
()3.1 Define, for $t\in[0,T]$, $\mathcal{F}_{2}(t)=\int_{0}^{t}F_{2}(t){\rm d}\tau.$ Set $v(x,t)=u(x,t)+\mathcal{F}_{2}(t)$. Then
$\widetilde{L}v(x,t)=\widetilde{L}u(x,t)+\mathcal{F}_{2}^{\prime}(t)=Lu(x,t)+F_{2}(t)\geq 0.$
On $\Gamma_{t}$, one has $v(x,t)\geq u(x,t)\geq G_{2}(t)$. By the Maximum Principle for $\widetilde{L}$ and $v$, we have $v(x,t)\geq G_{2}(t)$ on $U_{T}$. Thus, $u(x,t)\geq G_{2}(t)-\mathcal{F}_{2}(t)$ on $U_{T}$, which proves (3.3).
()3.1 Define, for $t\in[0,T]$, $\mathcal{F}(t)=\int_{0}^{t}F(t){\rm d}\tau.$ Applying (i) and (ii) to $F_{1}=F$ and $F_{2}=F$, respectively, we have
$u(x,t)\leq G_{1}(t)+\mathcal{F}(t)\leq G(t)+\mathcal{F}(t)\text{ and }u(x,t)\geq G_{2}(t)-\mathcal{F}(t)\geq-G(t)-\mathcal{F}(t),$
thus, obtain (3.4).
We consider Case ()3.1 now. Set $\widehat{L}w=w_{t}-\langle A,D^{2}w\rangle+\widetilde{b}\cdot\nabla w$, where
$\widetilde{b}(x,t)=u(x,t)B(x,t)+P^{\prime}(u(x,t))K(x,t)\nabla u(x,t).$
We have $Lu=\widehat{L}u$. By applying the results for Case ()3.1 to $\mathscr{L}:=\widehat{L}$, we obtain ()3.1, ()3.1 and ()3.1 again. ∎
The following is a simple consequence of Theorem 3.1. It shows that we can find preliminary estimates solutions of (2.9) for all time in terms of $f$, $g$ and $u_{0}$. These simple estimates will be improved much more later. Actually, they play the role of a starting point for those improvements.
Corollary 3.2.
Let $A(x,t)$ satisfy (3.1) with $T=\infty$ and let $u\in C(\overline{U_{\infty}})\cap C_{x,t}^{2,1}(U_{\infty})$ be a solution of (2.9). Assume $f^{+}(x,t)\leq F_{1}(t)$, resp., $f^{-}(x,t)\leq F_{2}(t)$ for all $(x,t)\in U_{\infty}$, where $F_{1}$, resp., $F_{2}$ is an integrable function in $C([0,\infty),[0,\infty))$, and $g(x,t)$ is bounded from above, resp., from below on $\Gamma\times(0,\infty)$. Denote
$m_{1}=\max\left\{\sup_{x\in U}u_{0}(x),\sup_{(x,t)\in\Gamma\times(0,\infty)}g(x,t)\right\}\text{ and }\widetilde{m}_{1}=\int_{0}^{\infty}F_{1}(\tau){\rm d}\tau,$ (3.5)
resp.,
$m_{2}=\min\left\{\inf_{x\in U}u_{0}(x),\inf_{(x,t)\in\Gamma\times(0,\infty)}g(x,t)\right\}\text{ and }\widetilde{m}_{2}=\int_{0}^{\infty}F_{2}(\tau){\rm d}\tau.$ (3.6)
Then one has
$u(\overline{U_{\infty}})\subset(-\infty,m_{1}+\widetilde{m}_{1}],\text{ resp.,}u(\overline{U_{\infty}})\subset[m_{2}-\widetilde{m}_{2},\infty).$ (3.7)
Proof.
For $(x,t)\in U_{\infty}$, we have
$Lu(x,t)=f(x,t)\leq f^{+}(x,t)\leq F_{1}(t),\text{ resp., }Lu(x,t)=f(x,t)\geq-f^{-}(x,t)\geq-F_{2}(t).$
For $t>0$, applying inequality (3.2), resp., (3.3) in Theorem 3.1 to $T:=2t$, we obtain
$\max_{\overline{U_{t}}}u\leq\max_{\Gamma_{t}}u+\int_{0}^{t}F_{1}(\tau){\rm d}\tau\leq\sup_{\Gamma_{\infty}}u+\int_{0}^{\infty}F_{1}(\tau){\rm d}\tau=m_{1}+\widetilde{m}_{1},$
resp.,
$\min_{\overline{U_{t}}}u\geq\min_{\Gamma_{t}}u-\int_{0}^{t}F_{2}(\tau){\rm d}\tau\geq\inf_{\Gamma_{\infty}}u-\int_{0}^{\infty}F_{2}(\tau){\rm d}\tau=m_{2}-\widetilde{m}_{2}.$
Thus, the statement (3.7) follows. ∎
4. The linear problem
Let $T$ be an extended number in $(0,\infty]$, and
$A:U_{T}\to\mathcal{M}^{n\times n}_{{\rm sym}}\text{ and }b:U_{T}\to\mathbb{R}^{n}\text{ be given functions.}$. (4.1)
Define the linear operator $\widetilde{L}$ by (2.10).
Assumption 4.1.
There are constants $c_{0}>0$ and $M_{1}>0$ such that (3.1) holds and
${\rm Tr}(A(x,t))\leq M_{1}\text{ for all }(x,t)\in U_{T}.$
Assumption 4.2.
There is a constant $M_{2}\geq 0$ so that
$|b(x,t)|\leq M_{2}\text{ for all }(x,t)\in U_{T}.$
We recall a Growth Lemma of Landis-type [12].
Lemma 4.3 (Growth Lemma [8, Lemma 4.2]).
Let Assumptions 4.1 and 4.2 hold for some number $T\in(0,\infty)$. Given a point $x_{0}\not\in\bar{U}$, let $r_{0}=r_{*}(x_{0})$ and $R=R_{*}(x_{0})$. Set
$\beta_{*}=\frac{M_{1}+M_{2}R}{2c_{0}},\ \beta=\max\left\{\beta_{*},\frac{R^{2}}{4c_{0}T}\right\},\ T_{*}=\frac{R^{2}}{4c_{0}\beta}\in(0,T],\ \eta_{*}=1-(r_{0}/R)^{2\beta}\in(0,1).$ (4.2)
Let
$w\in C(\overline{U_{T}})\cap C_{x,t}^{2,1}(U_{T}).$ (4.3)
If $\widetilde{L}w\leq 0$ on $U_{T}$ and $w\leq 0$ on $\Gamma\times[0,T]$, then one has
$\max_{x\in\bar{U}}w^{+}(x,T_{*})\leq\eta_{*}\max_{x\in\bar{U}}w^{+}(x,0).$
Notice in (4.2) that if
$4c_{0}\beta_{*}T=R^{2},$ (4.4)
then
$\beta=\beta_{*}=R^{2}/(4c_{0}T),\quad T_{*}=R^{2}/(4c_{0}\beta_{*})=T\text{ and }\eta_{*}=1-(r_{0}/R)^{2\beta_{*}}.$ (4.5)
Lemma 4.4 (Growth Lemma, inhomogeneous version).
We assume the same as in Lemma 4.3 up to (4.3) and also assume (4.4) holds. Suppose there is a function $F\in C([0,T],[0,\infty))$ such that $\widetilde{L}w(x,t)\leq F(t)$ for all $(x,t)\in U_{T}$. Then one has
$\max_{x\in\bar{U}}w^{+}(x,T)\leq\eta_{*}\max_{x\in\bar{U}}w^{+}(x,0)+\max_{(x,t)\in\Gamma\times[0,T]}w^{+}(x,t)+\int_{0}^{T}F(t){\rm d}t.$ (4.6)
Proof.
Denote $D=\int_{0}^{T}F(t){\rm d}t,$ $\delta=\max_{\Gamma\times[0,T]}w^{+}(x,t)$, $\Lambda=D+\delta$, and
$J(t)=\max_{x\in\bar{U}}w^{+}(x,t)\text{ for $t\in[0,T]$.}$
Define the function
$v(x,t)=w(x,t)-\delta-\int_{0}^{t}F(\tau){\rm d}\tau\text{ for }(x,t)\in\bar{U}\times[0,T].$
Then one has $v\in C(\overline{U_{T}})\cap C_{x,t}^{2,1}(U_{T})$ and
$\widetilde{L}v=\widetilde{L}w-F(t)\leq 0\text{ in }U_{T}.$ (4.7)
Moreover, for the boundary values of $v$, one has
$v(x,t)\leq w(x,t)-\delta\leq w^{+}(x,t)-\delta\leq 0\text{ on }\Gamma\times[0,T].$
Note that we now have the values of $\beta$, $T_{*}$ and $\eta_{*}$ in (4.5). By the previous Growth Lemma - Lemma 4.3 - applied to the solution $v$ of (4.7), we have
$\max_{x\in\bar{U}}v^{+}(x,T)\leq\eta_{*}\max_{x\in\bar{U}}v^{+}(x,0).$ (4.8)
On the one hand, $v(x,0)=w(x,0)-\delta\leq w(x,0)$, hence,
$\max_{x\in\bar{U}}v^{+}(x,0)\leq\max_{x\in\bar{U}}w^{+}(x,0)=J(0).$ (4.9)
On the other hand, $v(x,T)=w(x,T)-\delta-D=w(x,T)-\Lambda$, hence,
$w(x,T)=v(x,T)+\Lambda\leq\max_{x\in\bar{U}}v^{+}(x,T)+\Lambda.$
Thus,
$J(T)\leq\max_{x\in\bar{U}}v^{+}(x,T)+\Lambda.$ (4.10)
Combining (4.10) with (4.8) and (4.9), we obtain $J(T)\leq\eta_{*}J(0)+\Lambda$ which proves (4.6). ∎
We next apply Lemma 4.4 to consecutive intervals to obtain estimates, though still in very technical forms, for all time.
Proposition 4.5.
Let $T=\infty$ and assume the function $A(x,t)$ satisfies Assumption 4.1. Given a strictly increasing sequence $(T_{k})_{k=0}^{\infty}$ of nonnegative numbers, let $\tau_{k}=T_{k}-T_{k-1}$ for $k\geq 1$. For each $k\geq 1$, supposing
$|b(x,t)|\leq m_{k}\text{ on }U\times(T_{k-1},T_{k}]$ (4.11)
and letting $x_{k}$ be any point in $\mathbb{R}^{n}\setminus\bar{U}$, we set
$r_{k}=r_{*}(x_{k}),\ R_{k}=R_{*}(x_{k}),\ \beta_{k}=\frac{M_{1}+m_{k}R_{k}}{2c_{0}}\text{ and }\eta_{k}=1-(r_{k}/R_{k})^{2\beta_{k}}.$ (4.12)
Assume
$4c_{0}\beta_{k}\tau_{k}=R_{k}^{2}\text{ for all }k\geq 1$ (4.13)
Let $w\in C(\bar{U}\times[T_{0},\infty))\cap C_{x,t}^{2,1}(U\times(T_{0},\infty)$ and define
$J_{k}=\max_{x\in\bar{U}}w^{+}(x,T_{k})\text{ for $k\geq 0$.}$ (4.14)
Suppose there is a function $F\in C([T_{0},\infty),[0,\infty))$ such that
$\widetilde{L}w(x,t)\leq F(t)\text{ for all }(x,t)\in U\times(T_{0},\infty).$ (4.15)
Then one has, for all $k\geq 1$,
$\max_{(x,t)\in\bar{U}\times[T_{k-1},T_{k}]}w^{+}(x,t)\leq J_{k-1}+\Lambda_{k}$ (4.16)
and
$J_{k}\leq\sum_{m=0}^{k}\left[\left(\prod_{j=m+1}^{k}\eta_{j}\right)\Lambda_{m}\right],$ (4.17)
where
$\Lambda_{0}=J_{0}\text{ and }\Lambda_{k}=\max_{(x,t)\in\Gamma\times[T_{k-1},T_{k}]}w^{+}(x,t)+\int_{T_{k-1}}^{T_{k}}F(t){\rm d}t\text{ for $k\geq 1$. }$ (4.18)
Proof.
Thanks to (4.15), we can apply, for each $k\geq 1$, Theorem 3.1, case ()3.1, part ()3.1 to the interval $[0,\tau_{k}]$ in place of $[0,T]$, functions $u(t):=w(T_{k-1}+t)$, $F_{1}(t):=F(T_{k-1}+t)$ and select $t=\tau_{k}$ in (3.2). It results in
$\max_{\bar{U}\times[T_{k-1},T_{k}]}w$ $\leq\max_{(\bar{U}\times\{T_{k-1}\})\cup(\Gamma\times[T_{k-1},T_{k}])}w+\int_{T_{k-1}}^{T_{k}}F(\tau){\rm d}\tau$
$\leq\left(J_{k-1}+\max_{\Gamma\times[T_{k-1},T_{k}]}w^{+}\right)+\int_{T_{k-1}}^{T_{k}}F(\tau){\rm d}\tau\leq J_{k-1}+\Lambda_{k}$
which implies (4.16).
Let $k\geq 1$. We apply Lemma 4.4 to the interval $[T_{k-1},T_{k}]$ in place of $[0,T]$, with the same $M_{1}$, but, referring to the numbers in (4.11) and (4.12),
$M_{2}:=m_{k},\ x_{0}:=x_{k},\ r_{0}:=r_{k},\ R:=R_{k},\ \beta=\beta_{*}:=\beta_{k},\ \eta_{*}:=\eta_{k}.$
We obtain from (4.6) that
$J_{k}\leq\eta_{k}J_{k-1}+\Lambda_{k}.$
Iterating this estimate for $k-1$, $k-2$, etc., we have
$J_{k}$ $\leq\eta_{k}(\eta_{k-1}J_{k-2}+\Lambda_{k-1})+\Lambda_{k}=\eta_{k}\eta_{k-1}J_{k-2}+\eta_{k}\Lambda_{k-1}+\Lambda_{k}$
$\leq\eta_{k}\eta_{k-1}\eta_{k-2}J_{k-3}+\eta_{k}\eta_{k-1}\Lambda_{k-2}+\eta_{k}\Lambda_{k-1}+\Lambda_{k}$
$\leq\ldots\leq\eta_{k}\eta_{k-1}...\eta_{1}J_{0}+\sum_{m=1}^{k}\prod_{j=m+1}^{k}\eta_{j}\Lambda_{m},$
which implies (4.17). ∎
Below, we give more specific descriptions of the behavior of the solution $w(x,t)$, for large $t$ in the case $T=\infty$, of the equation $\widetilde{L}w(x,t)=f(x,t)$. For the rest of this section, $A$ and $b$ are fixed functions in (4.1) with $T=\infty$, and the linear operator $\widetilde{L}$ is correspondingly defined by (2.10) with $T=\infty$.
4.1. Case of bounded drifts
We first study the case when the drift term $b(x,t)$ is uniformly bounded in $x$ and $t$. We will apply Proposition 4.5 which raises the issue of estimating the series in (4.17). In some particular cases, we will use the following simple result.
Lemma 4.6.
Given a number $\eta\in(0,1)$ and a sequence $(\Lambda_{k})_{k=0}^{\infty}$ of nonnegative numbers. For $k\geq 0$, let $a_{k}=\sum_{j=0}^{k}\eta^{k-j}\Lambda_{j}$. Then
$\limsup_{k\to\infty}a_{k}\leq\frac{1}{1-\eta}\limsup_{k\to\infty}\Lambda_{k}.$ (4.19)
Proof.
If $(\Lambda_{k})_{k=0}^{\infty}$ is unbounded, then inequality (4.19) holds true. Now assume there is $\Lambda_{*}>0$ such that $\Lambda_{k}\leq\Lambda_{*}$ for all $k\geq 0$. For $m\geq 0$, let $\varepsilon_{m}=\sup\{\Lambda_{j}:j\geq m\}$. Then, for $k>m$, we have
$a_{k}$ $=\sum_{j=0}^{m}\eta^{k-j}\Lambda_{j}+\sum_{j=m+1}^{k}\eta^{k-j}\Lambda_{j}\leq\Lambda_{*}\sum_{j=0}^{m}\eta^{k-j}+\varepsilon_{m}\sum_{j=m+1}^{k}\eta^{k-j}$
$\leq\Lambda_{*}\eta^{k}\left(\sum_{j=0}^{m}\eta^{-j}\right)+\frac{\varepsilon_{m}}{1-\eta}.$
Passing $k\to\infty$ yields $\limsup_{k\to\infty}a_{k}\leq\varepsilon_{m}/(1-\eta).$ Then taking $m\to\infty$, we obtain (4.19). ∎
Our first explicit asymptotic estimates are the following.
Theorem 4.7.
Under Assumptions 4.1 and 4.2 both for $T=\infty$, let $w\in C(\overline{U_{\infty}})\cap C_{x,t}^{2,1}(U_{\infty})$. Given any point $x_{0}\not\in\bar{U}$, set the numbers $r_{0}=r_{*}(x_{0})$, $R=R_{*}(x_{0})$,
$\beta_{*}=\frac{1}{2c_{0}}(M_{1}+M_{2}R),\ T_{*}=\frac{R^{2}}{4c_{0}\beta_{*}}\text{ and }\eta_{*}=1-(r_{0}/R)^{2\beta_{*}}.$ (4.20)
  1. ()
    Suppose there is a functions $F_{1}\in C([0,\infty),[0,\infty))$ such that $\widetilde{L}w(x,t)\leq F_{1}(t)$ for all $(x,t)\in U_{\infty}$. Then
    $\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}w^{+}(x,t)\right]\leq\frac{2-\eta_{*}}{1-\eta_{*}}\left(\limsup_{t\to\infty}\left[\max_{x\in\Gamma}w^{+}(x,t)\right]+\limsup_{t\to\infty}\int_{t}^{t+T_{*}}F_{1}(\tau){\rm d}\tau\right).$ (4.21)
  2. ()
    Suppose there is a function $F\in C([0,\infty),[0,\infty))$ such that $|\widetilde{L}w(x,t)|\leq F(t)$ for all $(x,t)\in U_{\infty}$. Then
    $\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}|w(x,t)|\right]\leq\frac{2-\eta_{*}}{1-\eta_{*}}\left(\limsup_{t\to\infty}\left[\max_{x\in\Gamma}|w(x,t)|\right]+\limsup_{t\to\infty}\int_{t}^{t+T_{*}}F(\tau){\rm d}\tau\right).$ (4.22)
Proof.
()4.7 For $k\geq 0$, let $T_{k}=kT_{*}$ and $J_{k}=\max_{x\in\bar{U}}w^{+}(x,T_{k})$. Let $\Lambda_{0}=J_{0}$, and, for $k\geq 1$,
$\delta_{k}=\max_{\Gamma\times[T_{k-1},T_{k}]}w^{+},\quad D_{k}=\int_{T_{k-1}}^{T_{k}}F_{1}(t){\rm d}t,\quad\Lambda_{k}=D_{k}+\delta_{k}.$ (4.23)
We apply Proposition 4.5 to $x_{k}=x_{0}$, same $M_{1}$, and, referring to (4.11), $m_{k}=M_{2}$. Then $\tau_{k}=T_{k+1}-T_{k}=T_{*}$, the numbers in (4.12) are $r_{k}=r_{0}$, $R_{k}=R$, $\beta_{k}=\beta_{*}$, $\eta_{k}=\eta_{*}$ and condition (4.13) is met. It follows from (4.17) and Lemma 4.6 that
[…]
By (4.16) and (4.24),
[…]
One has from (4.23) that
[…]
Thus, by combining (4.25) with (4.26) we obtain the desired inequality (4.21).
()4.7 It suffices to consider the case when the right-hand side of (4.22) is a finite number which we denote by $C_{0}$. Applying part (i) to function $F_{1}:=F$, we obtain
$\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}w^{+}(x,t)\right]\leq\frac{2-\eta_{*}}{1-\eta_{*}}\left(\limsup_{t\to\infty}\left[\max_{\Gamma}w^{+}(x,t)\right]+\limsup_{t\to\infty}\int_{t}^{t+T_{*}}F(\tau){\rm d}\tau\right)\leq C_{0}.$ (4.27)
Note that $\widetilde{L}w\geq-F(t)$, hence $\widetilde{L}(-w)\leq F(t).$ Applying part (i) to functions $(-w)$ and $F_{1}(t):=F(t)$, we obtain, recalling that $(-w)^{+}=w^{-}$,
$\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}w^{-}(x,t)\right]\leq\frac{2-\eta_{*}}{1-\eta_{*}}\left(\limsup_{t\to\infty}\left[\max_{x\in\Gamma}w^{-}(x,t)\right]+\limsup_{t\to\infty}\int_{t}^{t+T_{*}}F(\tau){\rm d}\tau\right)\leq C_{0}.$ (4.28)
Observe that
$\max_{x\in\bar{U}}|w(x,t)|=\max_{x\in\bar{U}}\left[\max\{w^{+}(x,t),w^{-}(x,t)\}\right]\leq\max\left\{\max_{x\in\bar{U}}w^{+}(x,t),\max_{x\in\bar{U}}w^{-}(x,t)\right\}.$
Then one has
[…]
Combining this inequality (4.29) with the estimates (4.27) and (4.28), one obtains (4.22). ∎
Theorem 4.7 directly yields a result on the decay of $|w(x,t)|$.
Corollary 4.8.
Under Assumptions 4.1 and 4.2 both for $T=\infty$, let $w\in C(\overline{U_{\infty}})\cap C_{x,t}^{2,1}(U_{\infty})$ and assume there is a function $F\in C([0,\infty),[0,\infty))$ such that $|\widetilde{L}w(x,t)|\leq F(t)$ for all $(x,t)\in U_{\infty}$. If, in addition,
$\lim_{t\to\infty}\left(\max_{x\in\Gamma}|w(x,t)|\right)=0\text{ and }\lim_{t\to\infty}\int_{t}^{t+1}F(\tau){\rm d}\tau=0,$ (4.30)
then
$\lim_{t\to\infty}\left(\max_{x\in\bar{U}}|w(x,t)|\right)=0.$ (4.31)
Proof.
Let $T_{*}$ and $\eta_{*}$ be defined as in (4.20). We fix a positive integer $N\geq T_{*}$. Observe that
$\limsup_{t\to\infty}\int_{t}^{t+T_{*}}F(\tau){\rm d}\tau$ $\leq\limsup_{t\to\infty}\int_{t}^{t+N}F(\tau){\rm d}\tau=\limsup_{t\to\infty}\sum_{j=0}^{N-1}\int_{t+j}^{t+j+1}F(\tau){\rm d}\tau$ (4.32)
$\leq N\limsup_{t\to\infty}\int_{t}^{t+1}F(\tau){\rm d}\tau=0.$
Then (4.31) follows from (4.22), (4.30) and (4.32). ∎
Note that the second condition in (4.30) involves only the integral of $F$ on the time interval $[t,t+1]$. It does not require $F(t)\to 0$ as $t\to\infty$.
4.2. Case of unbounded drifts
Below, we study the case when $\sup_{x\in U}|b(x,t)|$ is unbounded as $t\to\infty$. Let $w\in C(\overline{U_{\infty}})\cap C_{x,t}^{2,1}(U_{\infty})$. The following technical conditions will be needed in our later proofs.
Assumption 4.9.
Assume there is $T^{*}\geq 0$ such that one has the following.
  1. ()
    There is a function $F\in C([T^{*},\infty),[0,\infty))$ such that
    $\widetilde{L}w(x,t)\leq F(t)\text{ for all }(x,t)\in U\times(T^{*},\infty).$ (4.33)
  2. ()
    There is a continuous function $z_{0}(t)>0$ on $[T^{*},\infty)$ increasing to infinity as $t\to\infty$ such that
    $|b(x,t)|\leq z_{0}(t)\text{ for all }(x,t)\in U\times(T^{*},\infty).$ (4.34)
  3. ()
    There is a decreasing, continuous function $\Lambda_{*}(t)\geq 0$ on $[T^{*},\infty)$ such that
    $\max_{(x,t)\in\Gamma\times[t,t+1]}w^{+}(x,t)+\int_{t}^{t+1}F(\tau){\rm d}\tau\leq\Lambda_{*}(t)\text{ for all }t\geq T^{*}.$ (4.35)
Referring to (4.35), we define a function $\bar{\Lambda}(t)$ on $[4T^{*},\infty)$ by
$\bar{\Lambda}(t)=\Lambda_{*}(t/4)\text{ for }t\geq 4T^{*}.$ (4.36)
Assumption 4.10.
Under Assumption 4.9, referring to the functions $z_{0}(t)$ in (4.34) and $\bar{\Lambda}(t)$ in (4.36), we assume
$\int_{T^{*}}^{\infty}e^{-\frac{d}{c_{0}}z_{0}(x)}{\rm d}x=\infty,$ (4.37)
and there are numbers $t_{*}\geq 4T^{*}$ and $\varepsilon_{*}>0$ such that, for any $\varepsilon\in(0,\varepsilon_{*})$, the function
$t\in[t_{*},\infty)\mapsto\bar{\Lambda}(t)\exp\left(e^{-\frac{d^{2}}{2c_{0}}-\varepsilon}\int_{t_{*}}^{t+1}e^{-\frac{d}{c_{0}}z_{0}(x)}{\rm d}x\right)\text{ is monotone.}$. (4.38)
For the upper bound of $w(x,t)$, we have the next theorem.
Theorem 4.11.
Under Assumption 4.1 with $T=\infty$ and Assumptions 4.9 and 4.10, if
$\limsup_{t\to\infty}\left(\bar{\Lambda}(t)e^{\frac{d}{c_{0}}z_{0}(t)}\right)=\ell\in[0,\infty),$ (4.39)
then $w(x,t)$ is bounded from above on $\overline{U_{\infty}}$ and
$\limsup_{t\to\infty}\left(\max_{x\in\bar{U}}w^{+}(x,t)\right)\leq e^{\frac{d^{2}}{2c_{0}}}\ell.$ (4.40)
Remark 4.12.
The following remarks are in order.
  1. ()
    If Assumption 4.9 holds for some $T^{*}\geq 0$, then, for any $\bar{T}>T^{*}$, Assumption 4.9 also holds with $T^{*}$ being replaced with $\bar{T}$.
  2. ()
    If Assumption 4.10 holds for some $T^{*}\geq 0$ and $t_{*}\geq 4T^{*}$, then, for any $\bar{T}\geq T^{*}$ and $\bar{t}\geq 4\bar{T}$, Assumption 4.10 also holds with $T^{*}$ and $t_{*}$ being replaced by $\bar{T}$ and $\bar{t}$ respectively.
  3. ()
    The limit superior (4.39) does not depend on the choice of $T^{*}$ and $t_{*}$ in Assumptions 4.9 and 4.10.
The proof of Theorem 4.11 involves constructing barrier functions with different centers. These centers will be selected to, at least, optimize the inequalities in (2.1). In fact, we have the following simple but useful fact.
Lemma 4.13.
For any $R>d$, there exists a point $y\not\in\bar{U}$ such that
$r_{*}(y)=R-d\text{ and }R_{*}(y)=R.$ (4.41)
Proof.
Let $\xi_{1},\xi_{2}\in\bar{U}$ such that $|\xi_{1}-\xi_{2}|=d>0$. First, let $y=y(t)\stackrel{{\scriptstyle\rm def}}{{=}}\xi_{1}+t(\xi_{2}-\xi_{1})$ with number $t>1$. We have $|y-\xi_{1}|=td$ and $|y-\xi_{2}|=(t-1)d$, which yield $|y-\xi_{1}|=d+|y-\xi_{2}|$.
Because $t>1$, one has $|y-\xi_{1}|>d$ which implies $y\not\in\bar{U}$.
For $z\in\bar{U}$, $|y-z|\leq|y-\xi_{2}|+|\xi_{2}-z|\leq|y-\xi_{2}|+d=|y-\xi_{1}|$. Hence, $|y-\xi_{1}|=R_{*}(y)$.
For $z\in\bar{U}$, $|y-z|\geq|y-\xi_{1}|-|\xi_{1}-z|\geq|y-\xi_{1}|-d=|y-\xi_{2}|$. Hence, $|y-\xi_{2}|=r_{*}(y)$.
Then the point $y=y(R/d)$ is not in $\bar{U}$ and it satisfies (4.41). ∎
Proof of Theorem 4.11.
Because of the continuity of $w$ on $\overline{U_{\infty}}$, it is bounded on $\overline{U_{T}}$ for any number $T\in(0,\infty)$. Thus, the estimate (4.40), once established, will imply that $w(x,t)$ is bounded from above on $\overline{U_{\infty}}$. Therefore, it suffices to prove (4.40) which we proceed in many steps below.
Step 1. Define the function
$Z(t)=-\frac{M_{1}+z_{0}(t)^{2}}{c_{0}}\ln\left(1-\frac{d}{z_{0}(t)}\right)\text{ for }t\geq T^{*}.$ (4.42)
By the Taylor expansion, we have
$-\ln(1-x)=x+\frac{1}{2}x^{2}+\mathcal{O}(x^{3})\text{ as }x\to 0.$
Consequently, as $t\to\infty$,
$Z(t)$ $=\frac{z_{0}(t)^{2}}{c_{0}}\left(\frac{d}{z_{0}(t)}+\frac{1}{2}\left(\frac{d}{z_{0}(t)}\right)^{2}+\mathcal{O}(1/z_{0}(t)^{3})\right)-\frac{M_{1}}{c_{0}}\ln\left(1-\frac{d}{z_{0}(t)}\right)$
$=\frac{d}{c_{0}}z_{0}(t)+\frac{d^{2}}{2c_{0}}+\mathcal{O}(1/z_{0}(t))+\mathcal{O}(1/z_{0}(t)).$ (4.43)
Let $\varepsilon_{0}$ be any number in the interval $(0,\varepsilon_{*})$. Thanks to (4.43), we can take a sufficiently large number $T_{0}>\max\{1/3,t_{*}\}$ such that
$z_{0}(T_{0})>\max\left\{d,\sqrt{M_{1}}\right\}\text{ and }Z(t)\leq\frac{d}{c_{0}}z_{0}(t)+\frac{d^{2}}{2c_{0}}+\varepsilon_{0}\text{ for all }t\geq T_{0}.$ (4.44)
For $t\geq 0$, define
$z_{*}(t)=z_{0}(T_{0}+t)\text{ and }z(t)=\frac{d}{c_{0}}z_{*}(t)+\frac{d^{2}}{2c_{0}}+\varepsilon_{0}=\frac{d}{c_{0}}z_{0}(T_{0}+t)+\frac{d^{2}}{2c_{0}}+\varepsilon_{0}.$ (4.45)
We then rewrite the inequality for $Z(t)$ in (4.44) as
$Z(T_{0}+t)\leq\frac{d}{c_{0}}z_{0}(T_{0}+t)+\frac{d^{2}}{2c_{0}}+\varepsilon_{0}=z(t)\text{ for all }t\geq 0.$ (4.46)
Note that the functions $z_{0}(t)$ and $z(t)$ are increasing and, for $t\geq 0$,
$z_{*}(t)\geq z_{0}(T_{0})>d.$ (4.47)
Step 2. For each integer $k\geq 1$, by (4.47) and the virtue of Lemma 4.13, we select a point $x_{k}\not\in\bar{U}$ so that
$R_{k}:=R_{*}(x_{k})=z_{*}(k),\quad r_{k}:=r_{*}(x_{k})=R_{k}-d,$ (4.48)
and define
$\widetilde{\beta}_{k}=\frac{M_{1}+R_{k}z_{*}(k)}{2c_{0}},\quad\tau_{k}=\frac{R_{k}^{2}}{4c_{0}\widetilde{\beta}_{k}}=\frac{R_{k}^{2}}{2(M_{1}+R_{k}z_{*}(k))}.$
For $k\geq 1$, set $T_{k}=T_{0}+\sum_{j=1}^{k}\tau_{j}$. We note from the choice (4.48) that
$\tau_{k}\leq\frac{R_{k}^{2}}{2R_{k}z_{*}(k)}=\frac{1}{2}.$ (4.49)
Also, by (4.44), the increase of $z_{*}(t)$ and (4.48), $M_{1}\leq z_{0}(T_{0})^{2}=z_{*}(0)^{2}\leq z_{*}(k)^{2}=R_{k}^{2}$, hence
$\tau_{k}=\frac{R_{k}^{2}}{2(M_{1}+R_{k}z_{*}(k))}\geq\frac{R_{k}^{2}}{2(R_{k}^{2}+R_{k}^{2})}=\frac{1}{4}.$
Therefore, for all $k\geq 0$,
$T_{0}+k/4\leq T_{k}\leq T_{0}+k/2.$ (4.50)
Step 3. We will apply Proposition 4.5. Let $k\geq 1$. Noticing from (4.50) that $[T_{k-1},T_{k}]\subset[T_{0},T_{0}+k]$, hence, we have on $U\times(T_{k-1},T_{k}]$
$|b(x,t)|\leq z_{0}(t)\leq z_{0}(T_{k})\leq z_{0}(T_{0}+k)=z_{*}(k).$
This prompts us to define $m_{k}=z_{*}(k)$ so that inequality (4.11) holds. Referring to definitions of the numbers $\beta_{k}$ and $\eta_{k}$ in Proposition 4.5, we have
$\beta_{k}=\frac{M_{1}+R_{k}m_{k}}{2c_{0}}=\frac{M_{1}+R_{k}z_{*}(k)}{2c_{0}}=\widetilde{\beta}_{k},$ (4.51)
$\eta_{k}=1-(r_{k}/R_{k})^{2\beta_{k}}=1-(r_{k}/R_{k})^{2\widetilde{\beta}_{k}}.$ (4.52)
Clearly, $4c_{0}\beta_{k}\tau_{k}=4c_{0}\widetilde{\beta}_{k}\tau_{k}=R_{k}^{2}$, hence condition (4.13) is met.
Step 4. For $k\geq 1$, we write $\eta_{k}$ in (4.52), with the use of (4.48) and (4.51), explicitly as
$\eta_{k}=1-\left(\frac{z_{*}(k)-d}{z_{*}(k)}\right)^{\frac{M_{1}+z_{*}^{2}(k)}{c_{0}}}=1-e^{-\tilde{z}(k)},$
where
$\tilde{z}(t)=-\frac{M_{1}+z_{*}(t)^{2}}{c_{0}}\ln\left(1-\frac{d}{z_{*}(t)}\right)\text{ for $t\geq 0$.}$
Note from (4.42) and (4.45) that $\tilde{z}(t)=Z(T_{0}+t)$. This identity and and property (4.46) imply $\tilde{z}(t)\leq z(t)$ for all $t\geq 0$. Therefore,
$\eta_{k}\leq 1-e^{-z(k)}\text{ for $k\geq 1$.}$ (4.53)
Step 5. Let $J_{k}$ and $\Lambda_{k}$ be defined by (4.14) and (4.18) respectively. For $k\geq 1$, by using the fact $\tau_{k}\leq 1$ in (4.49), we have
$\Lambda_{k}\leq\max_{\Gamma\times[T_{k-1},T_{k-1}+1]}w^{+}(x,t)+\int_{T_{k-1}}^{T_{k-1}+1}F(t){\rm d}t\leq\Lambda_{*}(T_{k-1}).$ (4.54)
Because $\Lambda_{*}(t)$ is decreasing and $03T_{0}-1>$ 0, we have for $k\geq 1$,
$\Lambda_{*}(T_{k-1})\leq\Lambda_{*}\left(T_{0}+\frac{k-1}{4}\right)=\Lambda_{*}\left(\frac{T_{0}+k+3T_{0}-1}{4}\right)\leq\Lambda_{*}\left(\frac{T_{0}+k}{4}\right).$ (4.55)
Define the function $\Lambda(t)$ by
$\Lambda(t)=\bar{\Lambda}(T_{0}+t)=\Lambda_{*}\left(\frac{T_{0}+t}{4}\right)\text{ for $t\geq 0$.}$
Then (4.54) and (4.55) yield
$\Lambda_{k}\leq\Lambda(k)\text{ for all }k\geq 1.$ (4.56)
Thanks to (4.39) and the fact $z_{0}(t)\to\infty$ as $t\to\infty$, one has $\Lambda(t)\to 0$ as $t\to\infty$. As a consequence of this and (4.56), one has
$\lim_{k\to\infty}\Lambda_{k}=0.$ (4.57)
Step 6. We now apply Proposition 4.5. By (4.16) and (4.57),
[…]
By inequality (4.17), we have, for $k\geq 1$,
$J_{k}\leq\sum_{m=0}^{k}(I_{k,m}\Lambda_{m}),\text{ where }I_{k,m}=\prod_{j=m+1}^{k}\eta_{j}.$ (4.59)
Let $k\geq 1$ and $0\leq m\leq k$. Using (4.53) and inequality $\ln(1+x)\leq x$, for $x\in(-1,\infty)$, we have
$\ln I_{k,m}\leq\sum_{j=m+1}^{k}\ln\left(1-e^{-z(j)}\right)\leq-\sum_{j=m+1}^{k}e^{-z(j)}.$
Because $z(j)$ is increasing in $j$, so is $-e^{-z(j)}$ and, hence, we can estimate
$-e^{-z(j)}\leq-\int_{j}^{j+1}e^{-z(x)}{\rm d}x.$
Thus, we obtain
$\ln I_{k,m}$ $\leq-\sum_{j=m+1}^{k}\int_{j}^{j+1}e^{-z(x)}{\rm d}x=-\int_{m+1}^{k+1}e^{-z(x)}{\rm d}x.$
Setting $E_{0}(t)=\exp\left(\int_{0}^{t}e^{-z(x)}{\rm d}x\right)\text{ for }t\geq 0$, we have $\ln I_{k,m}\leq E_{0}(m+1)-E_{0}(k+1)$. Therefore,
$I_{k,m}\leq\frac{E_{0}(m+1)}{E_{0}(k+1)},\text{ hence, }\sum_{m=0}^{k}I_{k,m}\Lambda_{m}\leq\frac{1}{E_{0}(k+1)}\sum_{m=0}^{k}\left[\Lambda_{m}E_{0}(m+1)\right].$
In the last sum, we consider $m=0$ separately and then use the relation (4.56) for $m\geq 1$. It yields that
$\sum_{m=0}^{k}I_{k,m}\Lambda_{m}$ $\leq\frac{\Lambda_{0}E_{0}(1)}{E_{0}(k+1)}+\frac{1}{E_{0}(k+1)}\sum_{m=1}^{k}\left[\Lambda(m)E_{0}(m+1)\right].$ (4.60)
By (4.45),
$E_{0}(t)=\exp\left(\int_{0}^{t}e^{-\frac{d}{c_{0}}z_{0}(T_{0}+x)}e^{-\frac{d^{2}}{2c_{0}}-\varepsilon_{0}}{\rm d}x\right)=\exp\left(e^{-\frac{d^{2}}{2c_{0}}-\varepsilon_{0}}\int_{T_{0}}^{T_{0}+t}e^{-\frac{d}{c_{0}}z_{0}(x)}{\rm d}x\right).$
Thanks to (4.37), we have $E_{0}(t)\to\infty$ as $t\to\infty$. Therefore, the first term on the right-hand side of (4.60) goes to zero as $k\to\infty$. Combining this with (4.59), we have
[…]
where
$\widetilde{J}_{k}=\frac{1}{E_{0}(k+1)}\sum_{m=1}^{k}\left[\Lambda(m)E_{0}(m+1)\right].$ (4.62)
Observe, for $t\geq 0$, that
$\Lambda(t)E_{0}(t+1)$ $=\Lambda(t)\exp\left(e^{-\frac{d^{2}}{2c_{0}}-\varepsilon_{0}}\int_{0}^{t+1}e^{-\frac{d}{c_{0}}z_{0}(T_{0}+x)}{\rm d}x\right)$
$=\bar{\Lambda}(T_{0}+t)\exp\left(e^{-\frac{d^{2}}{2c_{0}}-\varepsilon_{0}}\int_{T_{0}}^{T_{0}+t+1}e^{-\frac{d}{c_{0}}z_{0}(x)}{\rm d}x\right).$
Hence,
$\Lambda(t)E_{0}(t+1)=\frac{\bar{\Lambda}(T_{0}+t)\exp\left(e^{-\frac{d^{2}}{2c_{0}}-\varepsilon_{0}}\int_{t_{*}}^{T_{0}+t+1}e^{-\frac{d}{c_{0}}z_{0}(x)}{\rm d}x\right)}{\exp\left(e^{-\frac{d^{2}}{2c_{0}}-\varepsilon_{0}}\int_{t_{*}}^{T_{0}}e^{-\frac{d}{c_{0}}z_{0}(x)}{\rm d}x\right)}.$ (4.63)
Step 7. Regarding condition (4.38), we first consider the case the function
$q(t):=\bar{\Lambda}(t)\exp\left(e^{-\frac{d^{2}}{2c_{0}}-\varepsilon_{0}}\int_{t_{*}}^{t+1}e^{-\frac{d}{c_{0}}z_{0}(x)}{\rm d}x\right)\text{ is increasing on $[t_{*},\infty)$.}$ (4.64)
Denote the positive constant in the denominator in (4.63) by $E_{*}$. We rewrite (4.63) as
$\Lambda(t)E_{0}(t+1)=\frac{q(T_{0}+t)}{E_{*}}\text{ for }t\geq 0.$ (4.65)
Therefore, $\Lambda(t)E(t+1)$ is increasing. We use this fact to bound the summand in (4.60) from above by
$\Lambda(m)E_{0}(m+1)\leq\int_{m}^{m+1}E_{0}(s+1)\Lambda(s){\rm d}s.$ (4.66)
From (4.62) and (4.66), we obtain
$\widetilde{J}_{k}$ $\leq\frac{\sum_{m=1}^{k}\int_{m}^{m+1}E_{0}(s+1)\Lambda(s){\rm d}s}{E_{0}(k+1)}=\frac{\int_{1}^{k+1}E_{0}(s+1)\Lambda(s){\rm d}s}{E_{0}(k+1)}$ (4.67)
$=\int_{1}^{k+1}\exp\left(-\int_{s+1}^{k+1}e^{-z(x)}{\rm d}x\right)\Lambda(s){\rm d}s\leq y(k+1),$
where
$y(t)=\int_{0}^{t}\exp\left(-\int_{s+1}^{t}e^{-z(x)}{\rm d}x\right)\Lambda(s){\rm d}s.$ (4.68)
Rewrite
$y(t)=\int_{0}^{t}\exp\left(-\int_{s}^{t}e^{-z(x)}{\rm d}x\right)f(s){\rm d}s,\text{ where }f(s)=\exp\left(\int_{s}^{s+1}e^{-z(x)}{\rm d}x\right)\Lambda(s).$
We have, for $t>0$,
$y^{\prime}(t)=-e^{-z(t)}y(t)+f(t)=-h(t)\varphi^{-1}(y(t))+f(t),$ (4.69)
where $\varphi(x)=x$ and $h(t)=e^{-z(t)}$. Note from condition (4.37) that $\int_{0}^{\infty}h(\tau){\rm d}\tau=\infty.$ By applying [9, Lemma A.1] to equation (4.69), we obtain
$\limsup_{t\to\infty}y(t)$ $\leq\varphi\left(\limsup_{t\to\infty}\left[f(t)/h(t)\right]\right)=\limsup_{t\to\infty}\left[f(t)e^{z(t)}\right]$
$=\limsup_{t\to\infty}\left[\Lambda(t)e^{z(t)}\exp\left(\int_{t}^{t+1}e^{-z(x)}{\rm d}x\right)\right].$
Since $z(x)\to\infty$ as $x\to\infty$, one has $\int_{t}^{t+1}e^{-z(x)}{\rm d}x\to 0$ as $t\to\infty$. Consequently,
$\limsup_{t\to\infty}y(t)\leq\limsup_{t\to\infty}\left(\Lambda(t)e^{z(t)}\right)=e^{\frac{d^{2}}{2c_{0}}+\varepsilon_{0}}\limsup_{t\to\infty}\left(\bar{\Lambda}(T_{0}+t)e^{\frac{d}{c_{0}}z_{0}(T_{0}+t)}\right)=\ell e^{\frac{d^{2}}{2c_{0}}+\varepsilon_{0}}.$ (4.70)
It follows from (4.67) and (4.70) that
$\limsup_{k\to\infty}\widetilde{J}_{k}\leq\limsup_{k\in\mathbb{N},k\to\infty}y(k+1)\leq\limsup_{t\to\infty}y(t)\leq\ell e^{\frac{d^{2}}{2c_{0}}+\varepsilon_{0}}.$ (4.71)
Combining this with (4.61) gives
$\limsup_{k\to\infty}J_{k}\leq\ell e^{\frac{d^{2}}{2c_{0}}+\varepsilon_{0}}.$ (4.72)
Combining (4.58) with (4.72) yields
$\limsup_{t\to\infty}\left(\max_{x\in\bar{U}}w^{+}(x,t)\right)\leq\ell e^{\frac{d^{2}}{2c_{0}}+\varepsilon_{0}}.$
Passing $\varepsilon_{0}\to 0$, we obtain (4.40).
Step 8. Now, we consider, instead of (4.64), the case when the function $q(t)$ in (4.64) is decreasing. By (4.65), we have $\Lambda(t)E(t+1)$ is decreasing. With this property, we estimate, instead of (4.66),
$\Lambda(m)E_{0}(m+1)\leq\int_{m-1}^{m}E_{0}(s+1)\Lambda(s){\rm d}s.$
It implies
$\widetilde{J}_{k}$ $\leq\frac{\sum_{m=1}^{k}\int_{m-1}^{m}E_{0}(s+1)\Lambda(s){\rm d}s}{E_{0}(k+1)}=\frac{\int_{0}^{k}E_{0}(s+1)\Lambda(s){\rm d}s}{E_{0}(k+1)}=Y(k),$
where
$Y(t)=\frac{\int_{0}^{t}E_{0}(s+1)\Lambda(s){\rm d}s}{E_{0}(t+1)}.$
We rewrite
$Y(t)=\frac{\int_{0}^{t}\exp\left(-\int_{s+1}^{t}e^{-z(x)}{\rm d}x\right)\Lambda(s){\rm d}s}{\exp\left(\int_{t}^{t+1}e^{-z(x)}{\rm d}x\right)}=\frac{y(t)}{\exp\left(\int_{t}^{t+1}e^{-z(x)}{\rm d}x\right)}\leq y(t),$
where $y(t)$ is defined by (4.68). Then, together with (4.70),
$\limsup_{t\to\infty}Y(t)\leq\limsup_{t\to\infty}y(t)\leq\ell e^{\frac{d^{2}}{2c_{0}}+\varepsilon_{0}}.$
Thus,
$\limsup_{k\to\infty}\widetilde{J}_{k}\leq\limsup_{k\in\mathbb{N},k\to\infty}Y(k)\leq\limsup_{t\to\infty}Y(t)\leq\ell e^{\frac{d^{2}}{2c_{0}}+\varepsilon_{0}},$
that is, we obtain (4.71) again. Using the same arguments as in Step 7 after (4.71), we obtain (4.40) again. The proof is complete. ∎
We obtain an asymptotic estimate for $|w(x,t)|$ next.
Theorem 4.14.
Suppose we have the same assumptions as in Theorem 4.11 except that $|\widetilde{L}w|$ replaces $\widetilde{L}w$ in (4.33) and $|w(x,t)|$ replaces $w^{+}(x,t)$ in (4.35). Then one has
$\limsup_{t\to\infty}\left(\max_{x\in\bar{U}}|w(x,t)|\right)\leq e^{\frac{d^{2}}{2c_{0}}}\ell.$ (4.73)
Proof.
We apply Theorem 4.11 to $w$ and $(-w)$, and then use inequality (4.29) to obtain (4.73). We omit the details. ∎
Below, we have some examples for $z_{0}(t)$ and $\Lambda_{*}(t)$. Since $\Lambda_{*}(t)=\bar{\Lambda}(4t)$, it suffices to specify $\bar{\Lambda}$ instead of $\Lambda_{*}$.
Proposition 4.15.
Let $L>0$ and $\delta\geq 0$, and assume one of the following for sufficiently large $t>0$.
  1. ()
    For any number $\gamma>0$, let $z_{0}(t)=\gamma\ln\ln t$ and $\bar{\Lambda}(t)=L(\ln t)^{-\gamma\left(\frac{d}{c_{0}}+\delta\right)}$.
  2. ()
    For any positive number $\gamma<c_{0}/d$ and number $\alpha\in\mathbb{R}$, let $z_{0}(t)=\gamma\ln(t(\ln t)^{\alpha})$ and $\bar{\Lambda}(t)=L[t(\ln t)^{\alpha}]^{-\gamma\left(\frac{d}{c_{0}}+\delta\right)}$. In particular, if $\alpha=0$ then $z_{0}(t)=\gamma\ln t$ and $\bar{\Lambda}(t)=Lt^{-\gamma\left(\frac{d}{c_{0}}+\delta\right)}$.
  3. ()
    For any number $\alpha\in(-\infty,1)$, let $z_{0}(t)=\frac{c_{0}}{d}\ln(t(\ln t)^{\alpha})$ and $\bar{\Lambda}(t)=L[t(\ln t)^{\alpha}]^{-1-\frac{c_{0}\delta}{d}}$. In particular, if $\alpha=0$ then $z_{0}(t)=\frac{c_{0}}{d}\ln t$ and $\bar{\Lambda}(t)=Lt^{-1-\frac{c_{0}\delta}{d}}$.
Then the functions $z_{0}(t)$ and $\bar{\Lambda}(t)$ satisfy (4.37) and (4.38) for any $\varepsilon>0$ with some sufficiently large numbers $T^{*}$ and $t_{*}$. Moreover, the limit (4.39) holds with $\ell=L(1-{\rm sign}(\delta))$.
Proof.
One can verify that condition (4.37) is met in all the cases ()4.15, ()4.15, ()4.15 for a certain sufficiently large $T^{*}$. We check the remaining condition (4.38) for some large $t_{*}$ next. We construct the function $\bar{\Lambda}(t)$ in the following general form
$\bar{\Lambda}(t)=Le^{-\left(\frac{d}{c_{0}}+\delta\right)z_{0}(t)}.$ (4.74)
For $\varepsilon>0$, let
$E(t)=\exp\left(e^{-\frac{d^{2}}{2c_{0}}-\varepsilon}\int_{t_{*}}^{t+1}e^{-\frac{d}{c_{0}}z_{0}(x)}{\rm d}x\right)\text{ and }\mathcal{F}(t)=\bar{\Lambda}(t)E(t).$ (4.75)
It suffices to prove that $\mathcal{F}^{\prime}(t)$ has one sign for sufficiently large $t$. Taking the derivative of $\mathcal{F}(t)$ gives
$\mathcal{F}^{\prime}(t)$ $=-\bar{\Lambda}(t)\left(\frac{d}{c_{0}}+\delta\right)z_{0}^{\prime}(t)E(t)+\bar{\Lambda}(t)E(t)e^{-\frac{d^{2}}{2c_{0}}-\varepsilon}e^{-\frac{d}{c_{0}}z_{0}(t+1)}$ (4.76)
$=\bar{\Lambda}(t)E(t)\left\{-\left(\frac{d}{c_{0}}+\delta\right)z_{0}^{\prime}(t)+e^{-\frac{d^{2}}{2c_{0}}-\varepsilon}e^{-\frac{d}{c_{0}}z_{0}(t+1)}\right\}.$
Case ()4.15. For sufficiently large $t_{*}$, the calculations in (4.76) specifically yield
$\mathcal{F}^{\prime}(t)$ $=\bar{\Lambda}(t)E(t)\left\{-\left(\frac{d}{c_{0}}+\delta\right)\frac{\gamma}{t\ln t}+\frac{e^{-\frac{d^{2}}{2c_{0}}-\varepsilon}}{(\ln t)^{d\gamma/c_{0}}}\right\}.$
From this, one can easily see that $\mathcal{F}^{\prime}(t)>0$ for large $t$.
Case ()4.15. We have $z_{0}^{\prime}(t)=\frac{\gamma}{t}(1+\frac{\alpha}{\ln t}).$ Together with (4.76), this implies
$\mathcal{F}^{\prime}(t)$ $=\bar{\Lambda}(t)E(t)\left\{-\frac{\gamma}{t}\left(\frac{d}{c_{0}}+\delta\right)\left(1+\frac{\alpha}{\ln t}\right)+\frac{e^{-\frac{d^{2}}{2c_{0}}-\varepsilon}}{\left\{(t+1)(\ln(t+1))^{\alpha}\right\}^{d\gamma/c_{0}}}\right\}.$
Since $d\gamma/c_{0}<1$, we $\mathcal{F}^{\prime}(t)>0$ for large $t$.
Case ()4.15. Taking $\gamma=c_{0}/d$, we have, same as in ()4.15, that
$\mathcal{F}^{\prime}(t)=\bar{\Lambda}(t)E(t)\left\{-\frac{1}{t}\left(1+\gamma\delta\right)\left(1+\frac{\alpha}{\ln t}\right)+\frac{e^{-\frac{d^{2}}{2c_{0}}-\varepsilon}}{(t+1)(\ln(t+1))^{\alpha}}\right\}.$ (4.77)
If $\alpha\in(0,1)$, then the term $-(1+\gamma\delta)/t$ in (4.77) dominates as $t\to\infty$, hence $\mathcal{F}^{\prime}(t)<0$ for large $t$.
If $\alpha=0$, then (4.77) becomes
$\mathcal{F}^{\prime}(t)$ $=\bar{\Lambda}(t)E(t)\left\{-\frac{1}{t}\left(1+\gamma\delta\right)+\frac{e^{-\frac{d^{2}}{2c_{0}}-\varepsilon}}{t+1}\right\}.$
Because $e^{-\frac{d^{2}}{2c_{0}}-\varepsilon}<1\leq 1+\gamma\delta$, it follows that $\mathcal{F}^{\prime}(t)<0$ for large $t$.
If $\alpha<0$, then the last term in (4.77) dominates, for large $t$, and hence $\mathcal{F}^{\prime}(t)>0$. This completes the verification of condition (4.38) in Assumption 4.10.
Now, with $\bar{\Lambda}(t)$ in (4.74), it is obvious that the limit superior (4.39) is $\limsup_{t\to\infty}\left(Le^{-\delta z_{0}(t)}\right)$ which is $\ell=L$ when $\delta=0$, and is $\ell=0$ when $\delta>0$. This proves the last statement of this proposition. ∎
Below, we give more examples for the case $\ell=0$ in (4.39). These aim to remove the dependence of the functions $\bar{\Lambda}(t)$ in Proposition 4.15 on $\gamma$.
Proposition 4.16.
For any numbers $\alpha,\beta,\gamma,L>0$, let $z_{0}(t)=\gamma(\ln\ln t)^{\alpha}$ and $\bar{\Lambda}(t)=Lt^{-\beta}$. Then the functions $z_{0}(t)$ and $\bar{\Lambda}(t)$ satisfy (4.37) and (4.38) for any $\varepsilon>0$ with some sufficiently large numbers $T^{*}$ and $t_{*}$, and the limit (4.39) holds with $\ell=0$.
Proof.
Let $E(t)$ and $\mathcal{F}(t)$ be as in (4.75). Taking the derivative of $\mathcal{F}(t)$ gives
$\mathcal{F}^{\prime}(t)$ $=-\beta\bar{\Lambda}(t)t^{-1}E(t)+\bar{\Lambda}(t)E(t)e^{-\frac{d^{2}}{2c_{0}}-\varepsilon}e^{-\frac{d}{c_{0}}z_{0}(t+1)}=\bar{\Lambda}(t)E(t)\left\{-\frac{\beta}{t}+\frac{e^{-\frac{d^{2}}{2c_{0}}-\varepsilon}}{e^{\frac{\gamma d}{c_{0}}[\ln\ln(t+1)]^{\alpha}}}\right\}.$
For any $\varepsilon^{\prime}\in(0,1)$, one has, with sufficiently large $t$, $e^{\frac{\gamma d}{c_{0}}[\ln\ln(t+1)]^{\alpha}}\leq e^{\varepsilon^{\prime}\ln t}\leq t^{\varepsilon^{\prime}}$. Hence, $\mathcal{F}^{\prime}(t)>0$ for large $t$. Also, by choosing $\varepsilon^{\prime}<\beta$, one induces
$\bar{\Lambda}(t)e^{\frac{d}{c_{0}}z_{0}(t)}=Lt^{-\beta}e^{\frac{\gamma d}{c_{0}}[\ln\ln t]^{\alpha}}\to 0\text{ as }t\to\infty.$
This proves (4.39) with $\ell=0$. ∎
One can easily construct more complicated examples based on the above Propositions 4.15 and 4.16.
5. The nonlinear problem
We now return to the nonlinear equation (2.3) and study the initial boundary value problem (2.9). For a solution $u\in C(\overline{U_{\infty}})\cap C_{x,t}^{2,1}(U_{\infty})$ of (2.9), Theorem 3.1 and Corollary 3.2 already provide some rough lower and upper bounds. To improve these estimates, we need to take advantage of the inhomogeneous Growth Lemma, i.e., Lemma 4.4. It in turn requires a certain conversion of the nonlinear equation (2.3) to a suitable linear problem. We will do just that in subsection 5.1 below.
5.1. Reduction method
Recall the nonlinear operator $L$ defined by (2.8). To remove the quadratic terms of the gradient in $Lu$, we use a transformation of the Bernstein–Cole–Hopf type [2, 4, 11]. In this subsection 5.1, $T$ is an extended number in $(0,\infty]$, and $u$ is a function such that
$u\in C_{x,t}^{2,1}(U_{T})\text{ and $u(U_{T})\subset J$.}$
For any function $w\in C_{x,t}^{2,1}(U_{T})$, we define the function $\mathcal{L}w$ on $U_{T}$ by
$\mathcal{L}w=w_{t}-\langle A(x,t),D^{2}w\rangle+\widetilde{B}(x,t)\cdot\nabla w,\text{ where }\widetilde{B}(x,t)=B(x,t,u(x,t)).$ (5.1)
Clearly, $\mathcal{L}$ is a linear mapping from $C_{x,t}^{2,1}(U_{T})$ to the space of functions defined on $U_{T}$.
The following lemma is an extension of [8, Lemma 3.3] to inhomogeneous problems.
Lemma 5.1.
Below, for $\lambda\in\mathbb{R}$ and $C>0$, let $F_{\lambda,C}$ be any $C^{2}$-function on $J$ such that
$F^{\prime}_{\lambda,C}(s)=Ce^{\lambda P(s)}\text{ for any }s\in J.$ (5.2)
  1. ()
    Assume there is a constant $c_{1}\geq 0$ such that
    $\xi^{\rm T}K(x,t)\xi\geq-c_{1}|\xi|^{2}\text{ for all $(x,t)\in U_{T}$ and all$\xi\in\mathbb{R}^{n}$.}$ (5.3)
    For any numbers $\lambda\geq c_{1}/c_{0}$ and $C>0$, one has the function $w=F_{\lambda,C}(u)$ satisfies
    $\mathcal{L}w\leq F_{\lambda,C}^{\prime}(u)Lu\text{ on }U_{T}.$ (5.4)
  2. ()
    Assume there is a constant $c_{2}\geq 0$ such that
    $\xi^{\rm T}K(x,t)\xi\leq c_{2}|\xi|^{2}\text{ for all $(x,t)\in U_{T}$ and all$\xi\in\mathbb{R}^{n}$.}$ (5.5)
    For any numbers $\lambda\leq-c_{2}/c_{0}$ and $C>0$, one has the function $w=F_{\lambda,C}(u)$ satisfies
    $\mathcal{L}w\geq F_{\lambda,C}^{\prime}(u)Lu\text{ on }U_{T}.$ (5.6)
Proof.
Denoting $F=F_{\lambda,C}$, we have
$F^{\prime}(s)>0\text{ and }F^{\prime\prime}(s)=\lambda P^{\prime}(s)F^{\prime}(s)\text{ for any }s\in J.$
We obtain from [8, Eq. (3.18)] that
$\mathcal{L}w$ $=F^{\prime}(u)\{Lu-P^{\prime}(u)(\nabla u)^{\rm T}K(x,t)\nabla u\}-F^{\prime\prime}(u)(\nabla u)^{\rm T}A(x,t)\nabla u$ (5.7)
$=F^{\prime}(u)Lu-P^{\prime}(u)F^{\prime}(u)(\nabla u)^{\rm T}K(x,t)\nabla u-F^{\prime\prime}(u)(\nabla u)^{\rm T}A(x,t)\nabla u.$
()5.1 In this case with $\lambda\geq c_{1}/c_{0}\geq 0$, one has $F^{\prime\prime}(s)\geq 0$ and
$F^{\prime\prime}(s)\xi^{\rm T}A(x,t)\xi$ $\geq c_{0}F^{\prime\prime}(s)|\xi|^{2}=c_{0}\lambda P^{\prime}(s)F^{\prime}(s)|\xi|^{2}$ (5.8)
$\geq c_{1}P^{\prime}(s)F^{\prime}(s)|\xi|^{2}\geq-P^{\prime}(s)F^{\prime}(s)\xi^{\rm T}K(x,t)\xi$
for any $s\in J$ and $\xi\in\mathbb{R}^{n}$. The last inequality in (5.8) comes from (5.3). Applying (5.8) to $s=u(x,t)$, $\xi=\nabla u(x,t)$ and combining it with (5.7), we obtain $\mathcal{L}w\leq F^{\prime}(u)Lu$ which proves (5.4).
()5.1 In this case $\lambda\leq-c_{2}/c_{0}\leq 0$, one has $F^{\prime\prime}(s)\leq 0$ and
$F^{\prime\prime}(s)\xi^{\rm T}A(x,t)\xi$ $\leq c_{0}F^{\prime\prime}(s)|\xi|^{2}=c_{0}\lambda P^{\prime}(s)F^{\prime}(s)|\xi|^{2}$ (5.9)
$\leq-c_{2}P^{\prime}(s)F^{\prime}(s)|\xi|^{2}\leq-P^{\prime}(s)F^{\prime}(s)\xi^{\rm T}K(x,t)\xi$
for any $s\in J$ and $\xi\in\mathbb{R}^{n}$. Applying (5.9) to $s=u(x,t)$, $\xi=\nabla u(x,t)$ and combining it with (5.7), we obtain $\mathcal{L}w\geq F^{\prime}(u)Lu$ which proves (5.6). ∎
Note that the function $F_{\lambda,C}$ in Lemma 5.1 exists and is strictly increasing on $J$.
Example 5.2.
We recall the following examples from [8, Example 3.4].
  1. ()
    Case of isentropic, non-ideal gas flows. With the choice in (2.5), we can choose a particular function $F_{\lambda,C}$ as
[…]
  2. ()
    Case of ideal gas. With the choice in (2.6), we select $s_{0}=0$ in both cases of $J$. For $\lambda=0$, we clearly can choose $F_{\lambda,C}$ as
    $F_{\lambda}(s)=s\text{ for all }s\in J.$
    For $\lambda\neq 0$, we can choose $F_{\lambda,C}$ as
    $F_{\lambda}(s)=\frac{1}{\lambda}e^{\lambda s}\text{ for all }s\in J,\text{ or }F_{\lambda}(s)={\rm sign}(\lambda)e^{\lambda s}\text{ for all }s\in J.$ (5.11)
  3. ()
    Case of slightly compressible fluids. We use the choice in (2.4). When $\lambda\neq-1$, we can choose $F_{\lambda,C}$ as
    $F_{\lambda}(s)=\frac{1}{\lambda+1}s^{\lambda+1}\text{ for all }s>0,\text{ or }F_{\lambda}(s)={\rm sign}(\lambda+1)s^{\lambda+1}\text{ for all }s>0.$ (5.12)
    When $\lambda=-1$, we can choose $F_{\lambda,C}$ as $F_{\lambda}(s)=\ln(s)$ for all $s>0$.
5.2. Convergence of the solution
We finally utilize all the techniques developed previously to study the original inhomogeneous, nonlinear problem (2.9). For different types of function $B(x,t,s)$ in (2.8) and different scenarios of the forcing function and boundary data, we will establish the convergence of the solution, in the supremum norm in $x$, as time tends to infinity.
Assumption 5.3.
Let $A:U_{\infty}\to\mathcal{M}^{n\times n}_{{\rm sym}}$ and $K:U_{\infty}\to\mathcal{M}^{n\times n}$ be such that
  1. ()
    $A(x,t)$ satisfies Assumption 4.1 with $T=\infty$, and
  2. ()
    there are constants $c_{1}\geq 0$ and $c_{2}\geq 0$ such that
    $-c_{1}|\xi|^{2}\leq\xi^{\rm T}K(x,t)\xi\leq c_{2}|\xi|^{2}\text{ for all $(x,t)\in U_{\infty}$, all $\xi\in\mathbb{R}^{n}$.}$ (5.13)
For the remainder of this section, let Assumption 5.3 hold true. Let $u\in C(\overline{U_{\infty}})\cap C_{x,t}^{2,1}(U_{\infty})$ be a solution of (2.9). Let $\mathcal{L}$ be the linear mapping defined by (5.1) with $T=\infty$.
Thanks to condition (5.13) in Assumption 5.3, the matrix $K(x,t)$ satisfies both (5.3) and (5.5) with $T=\infty$.
Assumption 5.4.
Either $J=\mathbb{R}$ or one of the following cases holds true.
  1. ()
    If $J=[a,\infty)$ or $J=(-\infty,b]$, then $u(\overline{U_{\infty}})\subset J$.
  2. ()
    If $J=(a,\infty)$, then $u(\overline{U_{\infty}})\subset[m,\infty)$ for some $m>a$.
  3. ()
    If $J=(-\infty,b)$, then $u(\overline{U_{\infty}})\subset(-\infty,M]$ for some $M<b$.
  4. ()
    If $J$ is bounded, then $u(\overline{U_{\infty}})\subset[m,M]$ with $m,M\in J$.
Remark 5.5.
In the case of ideal gas flows (2.6) with $J=\mathbb{R}$, Assumption 5.4 trivially holds. In general, Assumption 5.4 is not meant to impose conditions on the solution $u(x,t)$ in $U_{\infty}$. Rather, when needed, it requires some conditions only on the forcing term $f(x,t)$, the boundary data $g(x,t)$, and the initial data $u_{0}(x)$ by the virtue of Corollary 3.2. For example, concerning the case ()5.4, resp. ()5.4, we only need to require that the numbers $m_{2}$ and $\widetilde{m}_{2}$ in (3.6), resp., $m_{1}$ and $\widetilde{m}_{1}$ in (3.5), satisfy $m_{2}-\widetilde{m}_{2}>a$, resp., $m_{1}+\widetilde{m}_{1}<b$.
Theorem 5.6 ($L_{t}^{1}$-forcing I).
In addition to Assumption 5.4, assume the following.
  1. ()
    There are constants $c_{B}>0$ and $\gamma_{0}>0$ such that
    $|B(x,t,s)|\leq c_{B}(1+|s|)^{\gamma_{0}}\text{ for all $(x,t)\in U_{\infty}$ and $s\in J$.}$ (5.14)
  2. ()
    There is a continuous function $F:[0,\infty)\to[0,\infty)$ such that
    $|f(x,t)|\leq F(t)\text{ for all $(x,t)\in U_{\infty}$.}$ (5.15)
  3. ()
    The function $F(t)$ in ()5.6 satisfies
    $\mu_{0}\stackrel{{\scriptstyle\rm def}}{{=}}\int_{0}^{\infty}F(\tau){\rm d}\tau<\infty.$ (5.16)
  4. ()
    There is $u_{*}\in J$ such that
    $\lim_{t\to\infty}\left(\max_{x\in\Gamma}|g(x,t)-u_{*}|\right)=0.$ (5.17)
Then one has
$\lim_{t\to\infty}\left(\max_{x\in\bar{U}}|u(x,t)-u_{*}|\right)=0.$ (5.18)
Proof.
The proof is divided into five steps.
Step 1. By (5.17), $u$ is bounded on $\Gamma\times(0,\infty)$. Thus, it is bounded on $\Gamma_{\infty}$ by some number $\mu_{1}>0$. By Theorem 3.1()3.1 and (5.16), we have $u$ is bounded on $\overline{U_{\infty}}$ by
$|u(x,t)|\leq\mu_{2}\stackrel{{\scriptstyle\rm def}}{{=}}\mu_{1}+\mu_{0}\text{ for all }(x,t)\in\overline{U_{\infty}}.$ (5.19)
By Assumption 5.4 and (5.19), we have
$u(\overline{U_{\infty}})\subset[m_{*},M_{*}]\text{ for some numbers $m_{*},M_{*}\in J$.}$ (5.20)
Let $\mu_{3}$ be a positive number such that
$|P(s)|\leq\mu_{3}\text{ for all }s\in[m_{*},M_{*}].$ (5.21)
Then
$|P(u(x,t))|\leq\mu_{3}\text{ for all }(x,t)\in\overline{U_{\infty}}.$ (5.22)
Thanks to properties (5.14) and (5.19), we have $\widetilde{B}(x,t)$ is bounded in $U_{\infty}$ by
$|\widetilde{B}(x,t)|\leq\mu_{4}\stackrel{{\scriptstyle\rm def}}{{=}}c_{B}(1+\mu_{2})^{\gamma_{0}}\text{ for all }(x,t)\in U_{\infty}.$ (5.23)
Step 2. Let
$\lambda_{1}>0\text{ and }\lambda_{2}<0\text{ such that }\lambda_{1}\geq c_{1}/c_{0}\text{ and }\lambda_{2}\leq-c_{2}/c_{0}.$
For $j=1,2$, let $F_{\lambda_{j}}$ be the function $F_{\lambda_{j},1}$ defined by (5.2), and define, thanks to (5.20),
$w_{j}=F_{\lambda_{j}}(u)\text{ on }\overline{U_{\infty}}.$
By Lemma 5.1 applied to $T=\infty$, $w:=w_{1}$ in part ()5.1 and $w:=w_{2}$ in part ()5.1, we have
$\mathcal{L}w_{1}\leq e^{\lambda_{1}P(u(x,t))}f(x,t)\text{ and }\mathcal{L}w_{2}\geq e^{\lambda_{2}P(u(x,t))}f(x,t)\text{ in }U_{\infty}.$ (5.24)
Denote $C_{1}=e^{\mu_{3}\max\{\lambda_{1},-\lambda_{2}\}}$. Then combining (5.22) with (5.15) gives
$e^{\lambda_{1}P(u(x,t))}|f(x,t)|\leq\mathcal{F}(t):=C_{1}F(t)\text{ and }e^{\lambda_{2}P(u(x,t))}|f(x,t)|\leq\mathcal{F}(t)\text{ in }U_{\infty}.$ (5.25)
For $j=1,2$, define $w_{*,j}=F_{\lambda_{j}}(u_{*})$ and $\bar{w}_{j}=w_{j}-w_{*,j}$ on $\overline{U_{\infty}}$. From (5.24) and (5.25) one has
$\mathcal{L}\bar{w}_{1}=\mathcal{L}w_{1}\leq\mathcal{F}(t)\text{ and }\mathcal{L}\bar{w}_{2}=\mathcal{L}w_{2}\geq-\mathcal{F}(t)\text{ in }U_{\infty}.$ (5.26)
Step 3. For $j=1,2$, by the local Lipschitz continuity of $F_{\lambda_{j}}(s)$ for $s\in J$ near $u_{*}$ and thanks to (5.17), there is a positive constant $C_{*}$ such that, for sufficiently large $t$ and all $x\in\Gamma$, one has
$|w_{j}(x,t)-w_{*,j}|=|F_{\lambda_{j}}(g(x,t))-F_{\lambda_{j}}(u_{*})|\leq C_{*}|g(x,t)-u_{*}|.$ (5.27)
Together with (5.17) again, this implies, for $j=1,2$,
$\lim_{t\to\infty}\left(\max_{x\in\Gamma}|\bar{w}_{j}(x,t)|\right)=0.$ (5.28)
Using the first inequality in (5.26), we can apply Theorem 4.7()4.7 to operator $\widetilde{L}:=\mathcal{L}$, function $w:=\bar{w}_{1}$, number $M_{2}:=\mu_{4}$ and function $F_{1}(t):=\mathcal{F}(t)$. We obtain from (4.21) that
$\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}\bar{w}_{1}^{+}(x,t)\right]\leq\frac{2-\eta_{*}}{1-\eta_{*}}\left(\limsup_{t\to\infty}\left[\max_{x\in\Gamma}\bar{w}_{1}^{+}(x,t)\right]+\limsup_{t\to\infty}\int_{t}^{t+T_{*}}\mathcal{F}(\tau){\rm d}\tau\right).$ (5.29)
For the last limit superior, it follows from condition (5.16) that
$\lim_{t\to\infty}\int_{t}^{t+T_{*}}\mathcal{F}(\tau){\rm d}\tau=C_{1}\lim_{t\to\infty}\int_{t}^{t+T_{*}}F(\tau){\rm d}\tau=0.$ (5.30)
Then, by combining (5.29) with the limits (5.28) and (5.30), one obtains
$\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}\bar{w}_{1}^{+}(x,t)\right]=0.$ (5.31)
Similarly, using the second inequality in (5.26), we can apply Theorem 4.7()4.7 to operator $\widetilde{L}=\mathcal{L}$, function $w:=-\bar{w}_{2}$, number $M_{2}=\mu_{4}$ and function $F_{1}(t):=\mathcal{F}(t)$. We obtain from (4.21) that
$\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}\bar{w}_{2}^{-}(x,t)\right]\leq\frac{2-\eta_{*}}{1-\eta_{*}}\left(\limsup_{t\to\infty}\left[\max_{x\in\Gamma}\bar{w}_{2}^{-}(x,t)\right]+\limsup_{t\to\infty}\int_{t}^{t+T_{*}}F(\tau){\rm d}\tau\right).$
Thus, thanks to (5.28) and (5.30) again,
$\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}\bar{w}_{2}^{-}(x,t)\right]=0.$ (5.32)
Step 4. Define $\bar{u}=u-u_{*}$ on $\overline{U_{\infty}}$. Denote $C_{2}=e^{\mu_{3}\min\{-\lambda_{1},\lambda_{2}\}}>0$. For $j=1,2$, we have from (5.2) and (5.21) that
$F_{\lambda_{j}}^{\prime}(s)=e^{\lambda_{j}P(s)}\geq C_{2}\text{ for }s\in[m_{*},M_{*}].$
By the Mean Value Theorem, one has, for $j=1,2$,
$C_{2}(s-u_{*})$ $\leq F_{\lambda_{j}}(s)-F_{\lambda_{j}}(u_{*})\text{ for }s\in[u_{*},M_{*}],$ (5.33)
$C_{2}(u_{*}-s)$ $\leq F_{\lambda_{j}}(u_{*})-F_{\lambda_{j}}(s)\text{ for }s\in[m_{*},u_{*}).$ (5.34)
When $j=1$, we have, based on (5.33), that
$(s-u_{*})^{+}\leq C_{2}^{-1}(F_{\lambda_{1}}(s)-F_{\lambda_{1}}(u_{*}))^{+}\text{ for all }s\in[m_{*},M_{*}].$ (5.35)
(For, (5.36) clearly holds true when $m_{*}\leq s\leq u_{*}$.) Similarly, when $j=2$, we have, based on (5.34), that
$(s-u_{*})^{-}\leq C_{2}^{-1}(F_{\lambda_{2}}(s)-F_{\lambda_{2}}(u_{*}))^{-}\text{ for all }s\in[m_{*},M_{*}].$ (5.36)
Applying (5.35) and (5.36) to $s=u(x,t)$ yields
$\bar{u}^{+}(x,t)\leq C_{2}^{-1}w_{1}^{+}(x,t)\text{ and }\bar{u}^{-}(x,t)\leq C_{2}^{-1}\bar{w}_{2}^{-}(x,t).$ (5.37)
Step 5. Thanks to inequalities (4.29), (5.37) and the limits (5.31), (5.32), we obtain
[…]
$\leq C_{2}^{-1}\max\left\{\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}\bar{w}_{1}^{+}(x,t)\right],\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}\bar{w}_{2}^{-}(x,t)\right]\right\}=0.$
Thus, we obtain (5.18). ∎
Next, we consider the case when $B(x,t,s)$ may not be uniformly bounded in $t\in[0,\infty)$. More specifically, the constant $c_{B}$ in (5.14) is replaced with an unbounded time-dependent function. For the sake of having concise statements below, we say that two nonnegative functions $z_{0}(t)$ and $\bar{\Lambda}(t)$, which are defined for sufficiently large $t$, satisfy Assumption 4.10 if the conditions from (4.37) to (4.38) are met for some numbers $T^{*}$, $t_{*}$ and $\varepsilon_{*}$.
Theorem 5.7 ($L_{t}^{1}$-forcing II).
In addition to Assumption 5.4, we assume the following.
  1. ()
    There is as continuous, increasing function $\bar{b}:[0,\infty)\to[0.\infty)$ which tends to infinity as $t\to\infty$, and a number $\gamma_{0}>0$ such that
    $|B(x,t,s)|\leq\bar{b}(t)(1+|s|)^{\gamma_{0}}\text{ for all }(x,t)\in U_{\infty}\text{ and }s\in J.$ (5.38)
  2. ()
    There is a continuous function $F:[0,\infty)\to[0,\infty)$ that satisfies (5.15) and (5.16).
  3. ()
    There are numbers $u_{*}\in J$, $T^{*}>0$ and a decreasing, continuous function $\widetilde{\Lambda}(t)\geq 0$ on $[T^{*},\infty)$ such that
    $\max_{(x,\tau)\in\Gamma\times[t,t+1]}\left|g(x,\tau)-u_{*}\right|+\int_{t}^{t+1}F(\tau){\rm d}\tau\leq\widetilde{\Lambda}(t)\text{ for all }t\geq T^{*}.$ (5.39)
  4. ()
    For any $\mu>0$, the functions $z_{0}(t):=\mu\bar{b}(t)$ and $\bar{\Lambda}(t):=\widetilde{\Lambda}(t/4)$ satisfy Assumption 4.10 and the limit (4.39) holds with $\ell:=\ell_{\mu}=0$.
Then one has (5.18).
Proof.
We follow the proof of Theorem 5.6 step by step.
Step 1. We proceed up to (5.22), and then replace (5.23) with the following, thanks to (5.38) and (5.19),
$|\widetilde{B}(x,t)|\leq\widehat{z}_{0}(t):=\mu_{6}\bar{b}(t)\text{ with }\mu_{6}=(1+\mu_{2})^{\gamma_{0}}.$ (5.40)
Step 2. Unchanged.
Step 3. We proceed up to (5.28). Let $\mu_{7}=\max\{C_{*},C_{1}\}$. For $j=1,2$, using (5.27), the definition of $\mathcal{F}(t)$ in (5.25), and the assumption (5.39) with function $\widetilde{\Lambda}(t)$, we can select a sufficiently large $T_{*}>0$ and have
$\max_{\Gamma\times[t,t+1]}|\bar{w}_{j}(x,t)|+\int_{t}^{t+1}\mathcal{F}(\tau){\rm d}\tau\leq\widehat{\Lambda}(t):=\mu_{7}\widetilde{\Lambda}(t)\text{ for all }t\geq T^{*}.$ (5.41)
For any numbers $\mu>0$ and $\gamma>0$, by assumption ()5.7, the functions $z_{0}(t):=\mu\bar{b}(t)$ and $\bar{\Lambda}(t):=\gamma\widetilde{\Lambda}(t/4)$ satisfy Assumption 4.10 and the limit (4.39) holds with
$\ell:=\ell_{\mu,\gamma}=\limsup_{t\to\infty}\left(\gamma\widetilde{\Lambda}(t/4)e^{\frac{d}{c_{0}}\mu z_{0}(t)}\right)=\gamma\limsup_{t\to\infty}\left(\widetilde{\Lambda}(t/4)e^{\frac{d}{c_{0}}\mu\bar{b}(t)}\right)=\gamma\ell_{\mu}=0.$
The last identity is, of course, due to the last statement in ()5.7.
(a) We consider the first inequality in (5.26) with $\bar{w}_{1}$. By the remark right after (5.41), the functions $z_{0}(t):=\widehat{z}_{0}(t)=\mu_{6}\bar{b}(t)$ and $\bar{\Lambda}(t):=\widehat{\Lambda}(t/4)=\mu_{7}\widetilde{\Lambda}(t/4)$ satisfy Assumption 4.10 and the limit (4.39) holds with $\ell:=\ell_{\mu_{6},\mu_{7}}=0$. We then apply Theorem 4.11 to $w:=\bar{w}_{1}$, $F(t):=\mathcal{F}(t)$, $z_{0}(t):=\widehat{z}_{0}(t)$ in (5.40), $\bar{\Lambda}(t):=\widehat{\Lambda}(t/4)$ in (5.41) and $\ell:=\ell_{\mu_{6},\mu_{7}}=0$. Thus, we have from (4.40) that
$\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}\bar{w}_{1}^{+}(x,t)\right]\leq e^{\frac{d^{2}}{2c_{0}}}\ell_{\mu_{6},\mu_{7}}=0,$
that is, we obtain inequality (5.31) again.
(b) Consider the second inequality in (5.26) with $\bar{w}_{2}$. Same as part (a), by applying Theorem 4.11 to $w:=-\bar{w}_{2}$, $F(t):=\mathcal{F}(t)$, $z_{0}(t):=\widehat{z}_{0}(t)$ and $\bar{\Lambda}(t):=\widehat{\Lambda}(t/4)$, we have
$\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}\bar{w}_{2}^{-}(x,t)\right]\leq e^{\frac{d^{2}}{2c_{0}}}\ell_{\mu_{6},\mu_{7}}=0.$
Thus, we obtain inequality (5.32) again.
Steps 4 and 5. Unchanged. At the end, we obtain the limit (5.18). ∎
Example 5.8.
Assume, for sufficiently large $t$,
$\bar{b}(t)=N_{1}(\ln\ln t)^{\alpha},\quad F(t)=N_{2}t^{-\beta_{0}},\quad\max_{x\in\Gamma}\left|g(x,t)-u_{*}\right|\leq N_{3}t^{-\beta_{1}},$
where $\alpha>0$, $\beta_{0}>1$, $\beta_{1}>0$ and $N_{1},N_{2},N_{3}>0$. Then the left-hand side of (5.39) is bounded by
$N_{3}t^{-\beta_{1}}+\frac{N_{2}}{\beta_{0}-1}t^{1-\beta_{0}}\leq\widetilde{\Lambda}(t):=N_{4}t^{-\beta_{2}},$
where $\beta_{2}=\min\{\beta_{1},\beta_{0}-1\}$ and $N_{4}>0$. By the virtue of Proposition 4.16, the condition ()5.7 in Theorem 5.7 is met.
Theorem 5.9 (Non-$L^{1}_{t}$ forcing I).
Consider
$J=(0,\infty)\text{ and $P(s)=\ln s$ for all $s\in J$.}$ (5.42)
Assume ()5.7 in Theorem 5.7 and ()5.6 in Theorem 5.6 and
$\int_{0}^{\infty}F(\tau){\rm d}\tau=\infty.$ (5.43)
Suppose there exist a number $u_{*}\in J$, a positive number $\lambda_{1}\geq c_{1}/c_{0}$, and two continuous functions $\mathcal{F}(t)$ and $\widetilde{\Lambda}(t)$ defined for sufficiently large $t$ such that the following are satisfied.
  1. ()
    The function $\mathcal{F}(t)$ is increasing and, for sufficiently large $t$,
    $\int_{0}^{t}F(\tau){\rm d}\tau\leq\mathcal{F}(t).$ (5.44)
  2. ()
    The function $\widetilde{\Lambda}(t)$ is decreasing and, for sufficiently large $t$,
    $\max_{(x,\tau)\in\Gamma\times[t,t+1]}|g(x,\tau)-u_{*}|+\int_{t}^{t+1}F(\tau)\mathcal{F}(\tau)^{\lambda_{1}}{\rm d}\tau\leq\widetilde{\Lambda}(t).$ (5.45)
  3. ()
    For any $\mu>0$, the functions $z_{0}(t):=\mu\bar{b}(t)\mathcal{F}(t)^{\gamma_{0}}$ and $\bar{\Lambda}(t):=\widetilde{\Lambda}(t/4)$ satisfy Assumption 4.10 and the limit (4.39) holds with $\ell:=\ell_{\mu}=0$.
If, additionally,
$u(\overline{U_{\infty}})\subset[m,\infty)\text{ for some number $m>0$,}$ (5.46)
then one has (5.18).
Proof.
We denote
$\Lambda_{g}(t)=\max_{(x,\tau)\in\Gamma\times[t,t+1]}|g(x,\tau)-u_{*}|.$ (5.47)
Step 1. Thanks to (5.43) and (5.44), one has
$\mathcal{F}(t)\to\infty\text{ as }t\to\infty.$ (5.48)
According to Condition ()5.9 with $\mu=1$, we have the limit
$\lim_{t\to\infty}\left\{\widetilde{\Lambda}(t/4)\exp\left(\frac{d}{c_{0}}\bar{b}(t)\mathcal{F}(t)^{\gamma_{0}}\right)\right\}=0.$ (5.49)
By the fact $\lim_{t\to\infty}\bar{b}(t)=\infty$ in Condition ()5.7 of Theorem 5.7 and and property (5.48) of $\mathcal{F}(t)$, we have from the limit (5.49) that $\widetilde{\Lambda}(t)\to 0$ as $t\to\infty$. This and (5.45) in turn imply
$\Lambda_{g}(t)\to 0\text{ as }t\to\infty.$ (5.50)
As a consequence of (5.50), one has
$u$ is bounded on $\Gamma_{\infty}$. (5.51)
By estimate (3.4) in Theorem 3.1()3.1, taking into account (5.51) and properties (5.44), (5.48), we can estimate, for $(x,t)\in\overline{U_{\infty}}$ with large $t$,
$|u(x,t)|\leq\mu_{1}\mathcal{F}(t),\text{ where $\mu_{1}$ is some positive number.}$. (5.52)
Together with (5.14), this implies, for $(x,t)\in U_{\infty}$ with large $t$,
$|\widetilde{B}(x,t)|\leq\widehat{z}_{0}(t):=\mu_{2}\bar{b}(t)\mathcal{F}(t)^{\gamma_{0}}\text{ for some number $\mu_{2}>0$.}$ (5.53)
Step 2. Let $\lambda_{2}<0$ such that $-1\neq\lambda_{2}\leq-c_{2}/c_{0}$. For $j=1,2$, define the function $w_{j}=F_{\lambda_{j}}(u):=\frac{1}{\lambda_{j}+1}u^{\lambda_{j}+1}$ on $\overline{U_{\infty}}$, see the first formula in (5.12) of Example 5.2()5.2. Let $\bar{u}=u-u_{*}$, and, for $j=1,2$, let $w_{*,j}=F_{\lambda_{j}}(u_{*})$ and define $\bar{w}_{j}=w_{j}-w_{*,j}$ on $\overline{U_{\infty}}$. Since property (5.50) implies (5.17), we obtain inequality (5.27) again, i.e., for all $x\in\Gamma$ and sufficiently large $t$,
$|\bar{w}_{j}(x,t)|\leq C_{*}|g(x,t)-u_{*}|\text{ for $j=1,2$.}$
Thus, for large $t$ and $j=1,2$,
$\max_{\Gamma\times[t,t+1]}|\bar{w}_{j}|\leq C_{*}\max_{\Gamma\times[t,t+1]}|g-u_{*}|=C_{*}\Lambda_{g}(t).$ (5.54)
Step 3. According to Lemma 5.1()5.1 with $\lambda=\lambda_{1}$ and $C=1$, one has, for $(x,t)\in U_{\infty}$, that
$\mathcal{L}w_{1}(x,t)\leq f(x,t)F_{\lambda_{1}}^{\prime}(u(x,t))=f(x,t)u(x,t)^{\lambda_{1}}.$ (5.55)
We use (5.15) to estimate $|f(x,t)|$ and use (5.52) to estimate $|u(x,t)|$ in (5.55). Then, for $(x,t)\in U_{\infty}$,
$\mathcal{L}\bar{w}_{1}=\mathcal{L}w_{1}\leq\mathcal{F}_{1}(t):=\mu_{3}F(t)\mathcal{F}^{\lambda_{1}}(t),\text{ where }\mu_{3}=\mu_{1}^{\lambda_{1}}.$ (5.56)
By (5.54) and the definition of $\mathcal{F}_{1}(t)$ in (5.56), we have, for large $t$,
$\max_{\Gamma\times[t,t+1]}|\bar{w}_{1}|+\int_{t}^{t+1}\mathcal{F}_{1}(\tau){\rm d}\tau\leq C_{*}\Lambda_{g}(t)+\mu_{3}\int_{t}^{t+1}F(\tau)\mathcal{F}(\tau)^{\lambda_{1}}{\rm d}\tau\leq\mu_{4}\widetilde{\Lambda}(t),$
where $\mu_{4}=\max\{C_{*},\mu_{3}\}$. Because of Condition ()5.9, by using the same argument as in Step 3 of the proof of Theorem 5.7 above, we can apply Theorem 4.11 to $w:=\bar{w}_{1}$, $F(t):=\mathcal{F}_{1}(t)$, $z_{0}(t):=\widehat{z}_{0}(t)$ in (5.53), $\bar{\Lambda}(t):=\mu_{4}\widetilde{\Lambda}(t/4)$ and
$\ell:=\ell_{\mu_{2},\mu_{4}}\stackrel{{\scriptstyle\rm def}}{{=}}\limsup_{t\to\infty}\left(\mu_{4}\widetilde{\Lambda}(t/4)e^{\frac{d}{c_{0}}\widehat{z}_{0}(t)}\right)=\mu_{4}\limsup_{t\to\infty}\left(\widetilde{\Lambda}(t/4)e^{\frac{d}{c_{0}}\mu_{2}\bar{b}(t)\mathcal{F}(t)^{\gamma_{0}}}\right)=\mu_{4}\ell_{\mu_{2}}=0.$ (5.57)
It follows from inequality (4.40) that
$\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}\bar{w}_{1}^{+}(x,t)\right]\leq\mu_{4}e^{\frac{d^{2}}{2c_{0}}}\ell=0.$ (5.58)
Since $\bar{w}_{1}\leq\bar{w}_{1}^{+}$, we obtain
$\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}\bar{w}_{1}(x,t)\right]\leq 0.$
Consequently,
$\limsup_{t\to\infty}\left[\max_{\bar{U}}w_{1}(x,t)\right]\leq w_{1,*},\text{ that is, }\limsup_{t\to\infty}\left[\max_{\bar{U}}F_{\lambda_{1}}(u(x,t))\right]\leq F_{\lambda_{1}}(u_{*}).$
From this we find that, see e.g. [8, Inequalities (4.60) and (4.61)],
$\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}u(x,t)\right]\leq u_{*}.$ (5.59)
Consequently,
$\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}\bar{u}(x,t)\right]\leq 0.$ (5.60)
Step 4. By (5.46), we have $u(x,t)\geq m>0$ for all $(x,t)\in\overline{U_{\infty}}$. According to Lemma 5.1()5.1 with $\lambda=\lambda_{2}<0$ and $C=1$, one has, for $(x,t)\in U_{\infty}$, that
$\mathcal{L}w_{2}\geq f(x,t)u^{\lambda_{2}}\geq-F(t)u^{\lambda_{2}}\geq-F(t)m^{\lambda_{2}}=-\mathcal{F}_{2}(t),\text{ where }\mathcal{F}_{2}(t)=F(t)m^{\lambda_{2}}.$
This implies
$\mathcal{L}\bar{w}_{2}=\mathcal{L}w_{2}\geq-\mathcal{F}_{2}(t).$
Consider sufficiently large $t$. By (5.48), we have $m^{\lambda_{2}}\leq\mathcal{F}(t)^{\lambda_{1}}$. Together with (5.54), this yields
$\max_{\Gamma\times[t,t+1]}|\bar{w}_{2}|+\int_{t}^{t+1}\mathcal{F}_{2}(\tau){\rm d}\tau\leq C_{*}\Lambda_{g}(t)+\int_{t}^{t+1}F(\tau)\mathcal{F}(\tau)^{\lambda_{1}}{\rm d}\tau\leq\mu_{5}\widetilde{\Lambda}(t),$
where $\mu_{5}=\max\{C_{*},1\}$. Applying Theorem 4.11 to $w:=-\bar{w}_{2}$, $F(t):=\mathcal{F}_{2}(t)$, $z_{0}(t):=\widehat{z}_{0}(t)$ in (5.53), $\bar{\Lambda}(t):=\mu_{5}\widetilde{\Lambda}(t/4)$ and, same as (5.57), $\ell:=\ell_{\mu_{2},\mu_{5}}=\mu_{5}\ell_{\mu_{2}}=0$, we obtain from (4.40) that
$\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}(-\bar{w}_{2}(x,t))^{+}\right]\leq e^{\frac{d^{2}}{2c_{0}}}\ell_{\mu_{2},\mu_{5}}=0.$
Hence,
$0\geq\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}(-\bar{w}_{2}(x,t))\right]=-\liminf_{t\to\infty}\left[\min_{x\in\bar{U}}\bar{w}_{2}(x,t)\right].$
This implies
$\liminf_{t\to\infty}\left[\min_{x\in\bar{U}}\bar{w}_{2}(x,t)\right]\geq 0.$
Therefore, one has, similar to (5.59),
$\liminf_{t\to\infty}\left[\min_{x\in\bar{U}}u(x,t)\right]\geq u_{*},$
and, consequently,
$\liminf_{t\to\infty}\left[\min_{x\in\bar{U}}\bar{u}(x,t)\right]\geq 0.$ (5.61)
Step 5. Let $\varepsilon>0$ be arbitrary. By (5.60) and (5.61), we have for sufficiently large $t$ that
$-\varepsilon\leq\min_{x\in\bar{U}}\bar{u}(x,t)\leq\bar{u}(x,t)\leq\max_{x\in\bar{U}}\bar{u}(x,t)\leq\varepsilon\text{ for all }x\in\bar{U},$
which implies $|\bar{u}(x,t)|\leq\varepsilon$ for all $x\in\bar{U}$. This proves the limit (5.18). ∎
Same as Remark 5.5, we comment again that condition (5.46) in Theorem 5.9 can be met thanks purely to the initial and boundary data. This comes from Corollary 3.2 with the condition $m_{2}>\widetilde{m}_{2}$. In that case, $m=m_{2}-\widetilde{m}_{2}>0$.
Example 5.10.
We give an example for the functions $\bar{b}(t)$, $F(t)$ and $\Lambda_{g}(t)$, see (5.47), so that Theorem 5.9 can apply. Let $\delta$ and $\beta$ be any numbers in the interval $(0,1)$. Assume, for sufficiently large $t$,
$\bar{b}(t)=N_{1}(\ln\ln t)^{\gamma_{1}},\quad F(t)=\frac{N_{2}}{t(\ln t)(\ln\ln t)^{\delta}},\quad\Lambda_{g}(t)\leq\frac{N_{3}}{t^{\beta}}$
for some positive numbers $\gamma_{1}$, $N_{1}$, $N_{2}$, $N_{3}$. Below, $N_{j}$, for $4\leq j\leq 7$, are appropriate positive numbers and $t$ is sufficiently large. We estimate
$\int_{0}^{t}F(\tau){\rm d}\tau\leq\mathcal{F}(t):=N_{4}(\ln\ln t)^{1-\delta}.$
Because $F(t)$ is decreasing and $\mathcal{F}(t)$ is increasing, one has
$\Lambda_{g}(t)+\int_{t}^{t+1}F(\tau)\mathcal{F}(\tau)^{\lambda_{1}}{\rm d}\tau\leq\Lambda_{g}(t)+F(t)\mathcal{F}(t+1)^{\lambda_{1}}\leq\widetilde{\Lambda}(t):=N_{5}t^{-\beta}.$
Let $\mu>0$ be arbitrary. Set
$z_{0}(t)=\mu\bar{b}(t)\mathcal{F}(t)^{\gamma_{0}}=N_{6}(\ln\ln t)^{\gamma_{2}},\text{ where }\gamma_{2}=\gamma_{1}+\gamma_{0}(1-\delta)>0.$
By Proposition 4.16, the functions $z_{0}(t)$ and $\bar{\Lambda}(t):=\widetilde{\Lambda}(t/4)=N_{7}t^{-\beta}$ satisfy Assumption 4.10 and the limit (4.39) holds with $\ell=0$.
The next result shows that the convergence in Theorem 5.9 can still hold true even when $u_{*}\not\in J$.
Theorem 5.11 (Non-$L^{1}_{t}$ forcing II).
Assume the same as in Theorem 5.9 with $u_{*}=0$, except for the last requirement (5.46). Then one has
$\lim_{t\to\infty}\left\{\max_{x\in\bar{U}}u(x,t)\right\}=0.$ (5.62)
Proof.
We follow the proof of Theorem 5.9. In this case $g\geq 0$, $u\geq 0$ and $u_{*}=0$. Although $P(s)$ is not defined at $s=u_{*}=0$, we have for $\lambda_{1}>0$ the transformation $F_{\lambda_{1}}(s)=\frac{s^{\lambda_{1}+1}}{\lambda_{1}+1}$ for $s>0$ can be extended to the function $\widehat{F}(s)=\frac{s^{\lambda_{1}+1}}{\lambda_{1}+1}$ for all $s\geq 0$. Note that $\widehat{F}^{\prime}(s)=s^{\lambda_{1}}$ is bounded near 00. In the proof of Theorem 5.9, we replace $F_{\lambda_{1}}$ with $\widehat{F}$ and proceed up to (5.58). In this case $\bar{w}_{1}=w_{1}=\widehat{F}(u)\geq 0$. Then inequality (5.58) reads as
$\limsup_{t\to\infty}\left[\frac{1}{\lambda_{1}+1}\max_{x\in\bar{U}}u^{\lambda_{1}+1}(x,t)\right]\leq 0.$
This implies, thanks to the fact $u\geq 0$, the desired limit (5.62). ∎
Remark 5.12.
The following final remarks are in order.
  1. ()
    Same as Theorem 5.11, one can consider (5.42) and $u_{*}=0$ for the $L_{t}^{1}$-forcing cases such as in Theorems 5.6 and 5.7.
  2. ()
    For more cases of $u(\overline{U_{\infty}})\subset\bar{J}$ and/or $u_{*}\in\partial J\setminus J$, see [8, Theorem 4.7].
  3. ()
    It is open that whether Theorem 5.9 can be established for isentropic fluid flows, see (2.5) and (2.6). The main obstacle is that the term $F_{\lambda_{1}}^{\prime}(u(x,t))$ in (5.55) can grow too large as $t\to\infty$, see (5.10) and (5.11), and, hence, Theorem 4.11 may not apply.
  4. ()
    In this paper, we studied only the Darcy law (1.2) for single-phase isothermal fluid flows. Other types of flows with nonlinear relations between $v$, $\rho$ and $\nabla p$ can also be considered, see e.g. [13, 1, 10, 3].
Data availability. No new data were created or analyzed in this study.
Funding. No funds were received for conducting this study.
Conflict of interest. There are no conflicts of interests.

### v2

@@
 $\frac{\partial\rho}{\partial t}=\langle A(x,t),D^{2}\rho\rangle-\frac{1}{\kappa\rho}(K_{0}(x,t)\nabla\rho)\cdot\nabla\rho+\rho B_{0}(x,t)\cdot\nabla\rho.$ (1.6)
-Equations (1.4), (1.5) and (1.6) fall into a larger class of equations which is formulated by (2.3).
+Equations (1.4), (1.5) and (1.6) fall into a larger class of equations which is formulated by (2.3) below.
 This nonlinear parabolic equation in non-divergence form will be the subject of our study.
 To the best of our knowledge, there are not many papers dealing with the long-time dynamics of non-divergence parabolic equations, especially the highly nonlinear ones such as (2.3).
-(However, see [14] for results for non-divergence linear elliptic problems with unbounded drift.) In our previous work [8], we essentially study (2.3) with zero forcing and, after shifting the solution by a constant, the homogeneous boundary condition.
+See [12, 13] for the general existence and regularity theory.
+See also [16] for results for non-divergence linear elliptic problems with unbounded drift.
+In our previous work [9], we essentially study (2.3) with zero forcing and, after shifting the solution by a constant, the homogeneous boundary condition.
 The purpose of the current paper is to obtain results for inhomogeneous problems.
@@
 For two matrices $A,B\in\mathcal{M}^{n\times n}$, their inner product $\langle A,B\rangle$ is the trace ${\rm Tr}(A^{\rm T}B)$.
-For a real-valued function $f(x)$ with $x=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}$, we denote by $D^{2}f$ the Hessian matrix of the second partial derivatives $(\partial^{2}f/\partial x_{i}\partial x_{j})_{i,j=1,\ldots,n}$.
+For a scalar function $f(x)$ with $x=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}$, we denote by $D^{2}f$ the Hessian matrix of the second partial derivatives $(\partial^{2}f/\partial x_{i}\partial x_{j})_{i,j=1,\ldots,n}$.
 For a real-valued function $f$, we recall that the positive and negative parts of $f$ are $f^{+}=\max\{0,f\}$ and $f^{-}=\max\{0,-f\}=(-f)^{+}$. Then one has
@@
 Then one has, for any $y\not\in\bar{U}$,
-$0<R_{*}(y)-d\leq r_{*}(y)<R_{*}(y)\leq r_{*}(y)+d.$ (2.1)
+$R_{*}(y)-d\leq r_{*}(y)<R_{*}(y)\leq r_{*}(y)+d.$ (2.1)
 We denote by $C_{x,t}^{2,1}(U\times I)$, for any interval $I$ of $\mathbb{R}$, the class of continuous functions $u(x,t)$ on $U\times I$ with continuous partial derivatives $\partial u/\partial t$, $\partial u/\partial x_{i}$, $\partial^{2}u/\partial x_{i}\partial x_{j}$ for $i,j=1.\ldots,n$.
@@
 $\frac{\partial u}{\partial t}-\langle A(x,t),D^{2}u\rangle+B(x,t,u)\cdot\nabla u+\mu(u)(K(x,t)\nabla u)\cdot\nabla u=f(x,t),$ (2.7)
-where $\mu$ is a continuous function from $J$ to $[0,\infty)$ or to $(-\infty,0]$.
-Indeed, in the former case, we just take $P$ such that $P^{\prime}=\mu$, while in the latter case, we take $P^{\prime}=-\mu$ and the new $K:=-K$.
+where $\mu$ is a continuous function from $J$ to $0,\infty)$ or to $(-\infty,0]$.
+Indeed, in the former case, we just take $P$ such that $P^{\prime}=\mu$, while in the latter case, we take $P^{\prime}=-\mu$ and the new K:=KK:=-K.
 Then equation (2.7) is of the form (2.3) with the function $P$ satisfying (2.2).
@@
 Thus, $u(x,t)\geq G_{2}(t)-\mathcal{F}_{2}(t)$ on $U_{T}$, which proves (3.3).
-()3.1 Define, for $t\in[0,T]$, $\mathcal{F}(t)=\int_{0}^{t}F(t){\rm d}\tau.$ Applying (i) and (ii) to $F_{1}=F$ and $F_{2}=F$, respectively, we have
+()3.1 Define, for $t\in[0,T]$, $\mathcal{F}(t)=\int_{0}^{t}F(t){\rm d}\tau.$ Applying (i) and ($i$) to $F_{1}=F$ and $F_{2}=F$, respectively, we have
 $u(x,t)\leq G_{1}(t)+\mathcal{F}(t)\leq G(t)+\mathcal{F}(t)\text{ and }u(x,t)\geq G_{2}(t)-\mathcal{F}(t)\geq-G(t)-\mathcal{F}(t),$
@@
 Then one has
-$u(\overline{U_{\infty}})\subset(-\infty,m_{1}+\widetilde{m}_{1}],\text{ resp.,}u(\overline{U_{\infty}})\subset[m_{2}-\widetilde{m}_{2},\infty).$ (3.7)
+[…]
 Proof.
@@
 Let $T$ be an extended number in $(0,\infty]$, and
-$A:U_{T}\to\mathcal{M}^{n\times n}_{{\rm sym}}\text{ and }b:U_{T}\to\mathbb{R}^{n}\text{ be given functions.}$. (4.1)
+$A:U_{T}\to\mathcal{M}^{n\times n}_{{\rm sym}}\text{ and }b:U_{T}\to\mathbb{R}^{n}\text{ be given functions.}$ (4.1)
 Define the linear operator $\widetilde{L}$ by (2.10).
@@
 We recall a Growth Lemma of Landis-type [14].
-Lemma 4.3 (Growth Lemma [8, Lemma 4.2]).
+Lemma 4.3 (Growth Lemma [9, Lemma 4.2]).
 Let Assumptions 4.1 and 4.2 hold for some number $T\in(0,\infty)$. Given a point $x_{0}\not\in\bar{U}$, let $r_{0}=r_{*}(x_{0})$ and $R=R_{*}(x_{0})$. Set
@@
 Assume
-$4c_{0}\beta_{k}\tau_{k}=R_{k}^{2}\text{ for all }k\geq 1$ (4.13)
-Let $w\in C(\bar{U}\times[T_{0},\infty))\cap C_{x,t}^{2,1}(U\times(T_{0},\infty)$ and define
+$4c_{0}\beta_{k}\tau_{k}=R_{k}^{2}\text{ for all }k\geq 1.$ (4.13)
+Let $w\in C(\bar{U}\times[T_{0},\infty))\cap C_{x,t}^{2,1}(U\times(T_{0},\infty))$ and define
 $J_{k}=\max_{x\in\bar{U}}w^{+}(x,T_{k})\text{ for $k\geq 0$.}$ (4.14)
@@
 Proof.
-Thanks to (4.15), we can apply, for each $k\geq 1$, Theorem 3.1, case ()3.1, part ()3.1 to the interval $[0,\tau_{k}]$ in place of $[0,T]$, functions $u(t):=w(T_{k-1}+t)$, $F_{1}(t):=F(T_{k-1}+t)$ and select $t=\tau_{k}$ in (3.2).
+Thanks to (4.15), we can apply, for each $k\geq 1$, Theorem 3.1, case ()3.1, part ()3.1 to the interval $0,\tau_{k}]$ in place of $[0,T]$, functions $u(t):=w(T_{k-1}+t)$, $F_{1}(t):=F(T_{k-1}+t)$ and select $t=\tau_{k}$ in (3.2).
 It results in
@@
 which implies (4.16).
-Let $k\geq 1$. We apply Lemma 4.4 to the interval $[T_{k-1},T_{k}]$ in place of $[0,T]$, with the same $M_{1}$, but, referring to the numbers in (4.11) and (4.12),
+Let $k\geq 1$. We apply Lemma 4.4 to the interval $T_{k-1},T_{k}]$ in place of $[0,T]$, with the same $M_{1}$, but, referring to the numbers in (4.11) and (4.12),
 $M_{2}:=m_{k},\ x_{0}:=x_{k},\ r_{0}:=r_{k},\ R:=R_{k},\ \beta=\beta_{*}:=\beta_{k},\ \eta_{*}:=\eta_{k}.$
@@
 $J_{k}\leq\eta_{k}J_{k-1}+\Lambda_{k}.$
-Iterating this estimate for $k-1$, $k-2$, etc., we have
+Iterating this estimate for k1k-1, k2k-2, etc., we have
 $J_{k}$ $\leq\eta_{k}(\eta_{k-1}J_{k-2}+\Lambda_{k-1})+\Lambda_{k}=\eta_{k}\eta_{k-1}J_{k-2}+\eta_{k}\Lambda_{k-1}+\Lambda_{k}$
@@
 It follows from (4.17) and Lemma 4.6 that
-[…]
+$\limsup_{k\to\infty}J_{k}\leq\limsup_{k\to\infty}\left(\sum_{m=0}^{k}\eta_{*}^{k-m}\Lambda_{m}\right)\leq\frac{1}{1-\eta_{*}}\limsup_{k\to\infty}\Lambda_{k}.$ (4.24)
 By (4.16) and (4.24),
-[…]
+$\limsup_{t\to\infty}\left(\max_{x\in\bar{U}}w^{+}(x,t)\right)\leq\limsup_{k\to\infty}J_{k-1}+\limsup_{k\to\infty}\Lambda_{k}\leq\frac{2-\eta_{*}}{1-\eta_{*}}\limsup_{k\to\infty}\Lambda_{k}.$ (4.25)
 One has from (4.23) that
-[…]
+$\limsup_{k\to\infty}\Lambda_{k}\leq\limsup_{k\to\infty}\delta_{k}+\limsup_{k\to\infty}D_{k}\leq\limsup_{t\to\infty}\left(\max_{x\in\Gamma}w^{+}(x,t)\right)+\limsup_{t\to\infty}\int_{t}^{t+T_{*}}F_{1}(\tau){\rm d}\tau.$ (4.26)
 Thus, by combining (4.25) with (4.26) we obtain the desired inequality (4.21).
@@
 $\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}w^{+}(x,t)\right]\leq\frac{2-\eta_{*}}{1-\eta_{*}}\left(\limsup_{t\to\infty}\left[\max_{\Gamma}w^{+}(x,t)\right]+\limsup_{t\to\infty}\int_{t}^{t+T_{*}}F(\tau){\rm d}\tau\right)\leq C_{0}.$ (4.27)
-Note that $\widetilde{L}w\geq-F(t)$, hence $\widetilde{L}(-w)\leq F(t).$ Applying part (i) to functions $(-w)$ and $F_{1}(t):=F(t)$, we obtain, recalling that $(-w)^{+}=w^{-}$,
+Note that $\widetilde{L}w\geq-F(t)$, hence $\widetilde{L}(-w)\leq F(t).$ Applying part (i) to functions (w)(-w) and $F_{1}(t):=F(t)$, we obtain, recalling that $(-w)^{+}=w^{-}$,
 $\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}w^{-}(x,t)\right]\leq\frac{2-\eta_{*}}{1-\eta_{*}}\left(\limsup_{t\to\infty}\left[\max_{x\in\Gamma}w^{-}(x,t)\right]+\limsup_{t\to\infty}\int_{t}^{t+T_{*}}F(\tau){\rm d}\tau\right)\leq C_{0}.$ (4.28)
@@
 Then one has
-[…]
+$\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}|w(x,t)|\right]\leq\max\left\{\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}w^{+}(x,t)\right],\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}w^{-}(x,t)\right]\right\}.$ (4.29)
 Combining this inequality (4.29) with the estimates (4.27) and (4.28), one obtains (4.22). ∎
@@
   2. ()
-    There is a continuous function $z_{0}(t)>0$ on $[T^{*},\infty)$ increasing to infinity as $t\to\infty$ such that
+    There is a continuous function $z_{0}(t)>0$ on $T^{*},\infty)$ increasing to infinity as $t\to\infty$ such that
     $|b(x,t)|\leq z_{0}(t)\text{ for all }(x,t)\in U\times(T^{*},\infty).$ (4.34)
   3. ()
-    There is a decreasing, continuous function $\Lambda_{*}(t)\geq 0$ on $[T^{*},\infty)$ such that
+    There is a decreasing, continuous function $\Lambda_{*}(t)\geq 0$ on $T^{*},\infty)$ such that
     $\max_{(x,t)\in\Gamma\times[t,t+1]}w^{+}(x,t)+\int_{t}^{t+1}F(\tau){\rm d}\tau\leq\Lambda_{*}(t)\text{ for all }t\geq T^{*}.$ (4.35)
-Referring to (4.35), we define a function $\bar{\Lambda}(t)$ on $[4T^{*},\infty)$ by
+Referring to (4.35), we define a function $\bar{\Lambda}(t)$ on $4T^{*},\infty)$ by
 $\bar{\Lambda}(t)=\Lambda_{*}(t/4)\text{ for }t\geq 4T^{*}.$ (4.36)
@@
 and there are numbers $t_{*}\geq 4T^{*}$ and $\varepsilon_{*}>0$ such that, for any $\varepsilon\in(0,\varepsilon_{*})$, the function
-$t\in[t_{*},\infty)\mapsto\bar{\Lambda}(t)\exp\left(e^{-\frac{d^{2}}{2c_{0}}-\varepsilon}\int_{t_{*}}^{t+1}e^{-\frac{d}{c_{0}}z_{0}(x)}{\rm d}x\right)\text{ is monotone.}$. (4.38)
+$t\in[t_{*},\infty)\mapsto\bar{\Lambda}(t)\exp\left(e^{-\frac{d^{2}}{2c_{0}}-\varepsilon}\int_{t_{*}}^{t+1}e^{-\frac{d}{c_{0}}z_{0}(x)}{\rm d}x\right)\text{ is monotone.}$ (4.38)
 For the upper bound of $w(x,t)$, we have the next theorem.
@@
 Step 1. Define the function
-$Z(t)=-\frac{M_{1}+z_{0}(t)^{2}}{c_{0}}\ln\left(1-\frac{d}{z_{0}(t)}\right)\text{ for }t\geq T^{*}.$ (4.42)
+[…]
 By the Taylor expansion, we have
-$-\ln(1-x)=x+\frac{1}{2}x^{2}+\mathcal{O}(x^{3})\text{ as }x\to 0.$
+$\ln(1-x)=x+\frac{1}{2}x^{2}+\mathcal{O}(x^{3})\text{ as }x\to 0.$
 Consequently, as $t\to\infty$,
@@
 $T_{0}+k/4\leq T_{k}\leq T_{0}+k/2.$ (4.50)
-Step 3. We will apply Proposition 4.5. Let $k\geq 1$. Noticing from (4.50) that $[T_{k-1},T_{k}]\subset[T_{0},T_{0}+k]$, hence, we have on $U\times(T_{k-1},T_{k}]$
+Step 3. We will apply Proposition 4.5. Let $k\geq 1$. Noticing from (4.50) that $T_{k-1},T_{k}]\subset[T_{0},T_{0}+k]$, hence, we have on $U\times(T_{k-1},T_{k}]$
 $|b(x,t)|\leq z_{0}(t)\leq z_{0}(T_{k})\leq z_{0}(T_{0}+k)=z_{*}(k).$
@@
 $\Lambda_{k}\leq\max_{\Gamma\times[T_{k-1},T_{k-1}+1]}w^{+}(x,t)+\int_{T_{k-1}}^{T_{k-1}+1}F(t){\rm d}t\leq\Lambda_{*}(T_{k-1}).$ (4.54)
-Because $\Lambda_{*}(t)$ is decreasing and $03T_{0}-1>$ 0, we have for $k\geq 1$,
+For $k\geq 1$, estimating $T_{k-1}$ by the first inequality in (4.50), and using the facts that $\Lambda_{*}(t)$ is decreasing and $3T_{0}-1>0$, we have
 $\Lambda_{*}(T_{k-1})\leq\Lambda_{*}\left(T_{0}+\frac{k-1}{4}\right)=\Lambda_{*}\left(\frac{T_{0}+k+3T_{0}-1}{4}\right)\leq\Lambda_{*}\left(\frac{T_{0}+k}{4}\right).$ (4.55)
@@
 Step 6. We now apply Proposition 4.5. By (4.16) and (4.57),
-[…]
+$\limsup_{t\to\infty}\left(\max_{x\in\bar{U}}w^{+}(x,t)\right)\leq\limsup_{k\to\infty}J_{k-1}+\limsup_{k\to\infty}\Lambda_{k}\leq\limsup_{k\to\infty}J_{k}.$ (4.58)
 By inequality (4.17), we have, for $k\geq 1$,
@@
 $\ln I_{k,m}\leq\sum_{j=m+1}^{k}\ln\left(1-e^{-z(j)}\right)\leq-\sum_{j=m+1}^{k}e^{-z(j)}.$
-Because $z(j)$ is increasing in $j$, so is $-e^{-z(j)}$ and, hence, we can estimate
-$-e^{-z(j)}\leq-\int_{j}^{j+1}e^{-z(x)}{\rm d}x.$
+Because $z(j)$ is increasing in $j$, so is $e^{-z(j)}$ and, hence, we can estimate
+$e^{-z(j)}\leq-\int_{j}^{j+1}e^{-z(x)}{\rm d}x.$
 Thus, we obtain
@@
 Thanks to (4.37), we have $E_{0}(t)\to\infty$ as $t\to\infty$. Therefore, the first term on the right-hand side of (4.60) goes to zero as $k\to\infty$. Combining this with (4.59), we have
-[…]
+$\limsup_{k\to\infty}J_{k}\leq\limsup_{k\to\infty}\sum_{m=0}^{k}I_{k,m}\Lambda_{m}\leq\limsup_{k\to\infty}\widetilde{J}_{k},$ (4.61)
 where
@@
 $y^{\prime}(t)=-e^{-z(t)}y(t)+f(t)=-h(t)\varphi^{-1}(y(t))+f(t),$ (4.69)
-where $\varphi(x)=x$ and $h(t)=e^{-z(t)}$. Note from condition (4.37) that $\int_{0}^{\infty}h(\tau){\rm d}\tau=\infty.$ By applying [9, Lemma A.1] to equation (4.69), we obtain
+where $\varphi(x)=x$ and $h(t)=e^{-z(t)}$. Note from condition (4.37) that $\int_{0}^{\infty}h(\tau){\rm d}\tau=\infty.$ By applying [8, Lemma A.1] to equation (4.69), we obtain
 $\limsup_{t\to\infty}y(t)$ $\leq\varphi\left(\limsup_{t\to\infty}\left[f(t)/h(t)\right]\right)=\limsup_{t\to\infty}\left[f(t)e^{z(t)}\right]$
@@
 Proof.
-We apply Theorem 4.11 to $w$ and $(-w)$, and then use inequality (4.29) to obtain (4.73). We omit the details. ∎
+We apply Theorem 4.11 to $w$ and (w)(-w), and then use inequality (4.29) to obtain (4.73). We omit the details. ∎
 Below, we have some examples for $z_{0}(t)$ and $\Lambda_{*}(t)$. Since $\Lambda_{*}(t)=\bar{\Lambda}(4t)$, it suffices to specify $\bar{\Lambda}$ instead of $\Lambda_{*}$.
@@
 $\mathcal{F}^{\prime}(t)=\bar{\Lambda}(t)E(t)\left\{-\frac{1}{t}\left(1+\gamma\delta\right)\left(1+\frac{\alpha}{\ln t}\right)+\frac{e^{-\frac{d^{2}}{2c_{0}}-\varepsilon}}{(t+1)(\ln(t+1))^{\alpha}}\right\}.$ (4.77)
-If $\alpha\in(0,1)$, then the term $-(1+\gamma\delta)/t$ in (4.77) dominates as $t\to\infty$, hence $\mathcal{F}^{\prime}(t)<0$ for large $t$.
+If $\alpha\in(0,1)$, then the term $(1+\gamma\delta)/t$ in (4.77) dominates as $t\to\infty$, hence $\mathcal{F}^{\prime}(t)<0$ for large $t$.
 If $\alpha=0$, then (4.77) becomes
@@
 Clearly, $\mathcal{L}$ is a linear mapping from $C_{x,t}^{2,1}(U_{T})$ to the space of functions defined on $U_{T}$.
-The following lemma is an extension of [8, Lemma 3.3] to inhomogeneous problems.
+The following lemma is an extension of [9, Lemma 3.3] to inhomogeneous problems.
 Lemma 5.1.
@@
 $F^{\prime}(s)>0\text{ and }F^{\prime\prime}(s)=\lambda P^{\prime}(s)F^{\prime}(s)\text{ for any }s\in J.$
-We obtain from [8, Eq. (3.18)] that
+We obtain from [9, Eq. (3.18)] that
 $\mathcal{L}w$ $=F^{\prime}(u)\{Lu-P^{\prime}(u)(\nabla u)^{\rm T}K(x,t)\nabla u\}-F^{\prime\prime}(u)(\nabla u)^{\rm T}A(x,t)\nabla u$ (5.7)
@@
 Example 5.2.
-We recall the following examples from [8, Example 3.4].
+We recall the following examples from [9, Example 3.4].
   1. ()
     Case of isentropic, non-ideal gas flows. With the choice in (2.5), we can choose a particular function $F_{\lambda,C}$ as
-[…]
+    $F_{\lambda}(s)=\int_{0}^{s}e^{\lambda z^{\gamma}}dz\text{ for all }s\geq 0.$ (5.10)
   2. ()
@@
     For $\lambda\neq 0$, we can choose $F_{\lambda,C}$ as
-    $F_{\lambda}(s)=\frac{1}{\lambda}e^{\lambda s}\text{ for all }s\in J,\text{ or }F_{\lambda}(s)={\rm sign}(\lambda)e^{\lambda s}\text{ for all }s\in J.$ (5.11)
+[…]
   3. ()
     Case of slightly compressible fluids. We use the choice in (2.4). When $\lambda\neq-1$, we can choose $F_{\lambda,C}$ as
-    $F_{\lambda}(s)=\frac{1}{\lambda+1}s^{\lambda+1}\text{ for all }s>0,\text{ or }F_{\lambda}(s)={\rm sign}(\lambda+1)s^{\lambda+1}\text{ for all }s>0.$ (5.12)
+[…]
     When $\lambda=-1$, we can choose $F_{\lambda,C}$ as $F_{\lambda}(s)=\ln(s)$ for all $s>0$.
@@
     there are constants $c_{1}\geq 0$ and $c_{2}\geq 0$ such that
-    $-c_{1}|\xi|^{2}\leq\xi^{\rm T}K(x,t)\xi\leq c_{2}|\xi|^{2}\text{ for all $(x,t)\in U_{\infty}$, all $\xi\in\mathbb{R}^{n}$.}$ (5.13)
+    $c_{1}|\xi|^{2}\leq\xi^{\rm T}K(x,t)\xi\leq c_{2}|\xi|^{2}\text{ for all $(x,t)\in U_{\infty}$, all $\xi\in\mathbb{R}^{n}$.}$ (5.13)
 For the remainder of this section, let Assumption 5.3 hold true.
@@
 By Assumption 5.4 and (5.19), we have
-$u(\overline{U_{\infty}})\subset[m_{*},M_{*}]\text{ for some numbers $m_{*},M_{*}\in J$.}$ (5.20)
+[…]
 Let $\mu_{3}$ be a positive number such that
@@
 $\lim_{t\to\infty}\left(\max_{x\in\Gamma}|\bar{w}_{j}(x,t)|\right)=0.$ (5.28)
-Using the first inequality in (5.26), we can apply Theorem 4.7()4.7 to operator $\widetilde{L}:=\mathcal{L}$, function $w:=\bar{w}_{1}$, number $M_{2}:=\mu_{4}$ and function $F_{1}(t):=\mathcal{F}(t)$.
+Using the first inequality in (5.26), we can apply Theorem 4.7()4.7 to operator $\widetilde{L}$$\mathcal{L}$, function $w:=\bar{w}_{1}$, number $M_{2}:=\mu_{4}$ and function $F_{1}(t):=\mathcal{F}(t)$.
 We obtain from (4.21) that
@@
 $\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}\bar{w}_{1}^{+}(x,t)\right]=0.$ (5.31)
-Similarly, using the second inequality in (5.26), we can apply Theorem 4.7()4.7 to operator $\widetilde{L}=\mathcal{L}$, function $w:=-\bar{w}_{2}$, number $M_{2}=\mu_{4}$ and function $F_{1}(t):=\mathcal{F}(t)$.
+Similarly, using the second inequality in (5.26), we can apply Theorem 4.7()4.7 to operator $\widetilde{L}$$\mathcal{L}$, function $w:=-\bar{w}_{2}$, number $M_{2}=\mu_{4}$ and function $F_{1}(t):=\mathcal{F}(t)$.
 We obtain from (4.21) that
@@
 Step 5. Thanks to inequalities (4.29), (5.37) and the limits (5.31), (5.32), we obtain
-[…]
+$\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}|\bar{u}(x,t)|\right]\leq\max\left\{\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}\bar{u}^{+}(x,t)\right],\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}\bar{u}^{-}(x,t)\right]\right\}$
 $\leq C_{2}^{-1}\max\left\{\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}\bar{w}_{1}^{+}(x,t)\right],\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}\bar{w}_{2}^{-}(x,t)\right]\right\}=0.$
@@
   3. ()
-    There are numbers $u_{*}\in J$, $T^{*}>0$ and a decreasing, continuous function $\widetilde{\Lambda}(t)\geq 0$ on $[T^{*},\infty)$ such that
+    There are numbers $u_{*}\in J$, $T^{*}>0$ and a decreasing, continuous function $\widetilde{\Lambda}(t)\geq 0$ on $T^{*},\infty)$ such that
     $\max_{(x,\tau)\in\Gamma\times[t,t+1]}\left|g(x,\tau)-u_{*}\right|+\int_{t}^{t+1}F(\tau){\rm d}\tau\leq\widetilde{\Lambda}(t)\text{ for all }t\geq T^{*}.$ (5.39)
@@
 $\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}\bar{w}_{1}^{+}(x,t)\right]\leq e^{\frac{d^{2}}{2c_{0}}}\ell_{\mu_{6},\mu_{7}}=0,$
-that is, we obtain inequality (5.31) again.
+that is, we obtain (5.31) again.
 (b) Consider the second inequality in (5.26) with $\bar{w}_{2}$.
@@
 $\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}\bar{w}_{2}^{-}(x,t)\right]\leq e^{\frac{d^{2}}{2c_{0}}}\ell_{\mu_{6},\mu_{7}}=0.$
-Thus, we obtain inequality (5.32) again.
+Thus, we obtain (5.32) again.
 Steps 4 and 5. Unchanged. At the end, we obtain the limit (5.18). ∎
@@
 If, additionally,
-$u(\overline{U_{\infty}})\subset[m,\infty)\text{ for some number $m>0$,}$ (5.46)
+[…]
 then one has (5.18).
@@
 By estimate (3.4) in Theorem 3.1()3.1, taking into account (5.51) and properties (5.44), (5.48), we can estimate, for $(x,t)\in\overline{U_{\infty}}$ with large $t$,
-$|u(x,t)|\leq\mu_{1}\mathcal{F}(t),\text{ where $\mu_{1}$ is some positive number.}$. (5.52)
+$|u(x,t)|\leq\mu_{1}\mathcal{F}(t),\text{ where $\mu_{1}$ is some positive number.}$ (5.52)
 Together with (5.14), this implies, for $(x,t)\in U_{\infty}$ with large $t$,
@@
 Step 2.
-Let $\lambda_{2}<0$ such that $-1\neq\lambda_{2}\leq-c_{2}/c_{0}$.
+Let $\lambda_{2}<0$ such that $1\neq\lambda_{2}\leq-c_{2}/c_{0}$.
 For $j=1,2$, define the function $w_{j}=F_{\lambda_{j}}(u):=\frac{1}{\lambda_{j}+1}u^{\lambda_{j}+1}$ on $\overline{U_{\infty}}$, see the first formula in (5.12) of Example 5.2()5.2.
@@
 $\mathcal{L}\bar{w}_{1}=\mathcal{L}w_{1}\leq\mathcal{F}_{1}(t):=\mu_{3}F(t)\mathcal{F}^{\lambda_{1}}(t),\text{ where }\mu_{3}=\mu_{1}^{\lambda_{1}}.$ (5.56)
-By (5.54) and the definition of $\mathcal{F}_{1}(t)$ in (5.56), we have, for large $t$,
+By (5.54), the definition of $\mathcal{F}_{1}(t)$ in (5.56), and inequality (5.45) we have, for large $t$,
 $\max_{\Gamma\times[t,t+1]}|\bar{w}_{1}|+\int_{t}^{t+1}\mathcal{F}_{1}(\tau){\rm d}\tau\leq C_{*}\Lambda_{g}(t)+\mu_{3}\int_{t}^{t+1}F(\tau)\mathcal{F}(\tau)^{\lambda_{1}}{\rm d}\tau\leq\mu_{4}\widetilde{\Lambda}(t),$
@@
 $\limsup_{t\to\infty}\left[\max_{\bar{U}}w_{1}(x,t)\right]\leq w_{1,*},\text{ that is, }\limsup_{t\to\infty}\left[\max_{\bar{U}}F_{\lambda_{1}}(u(x,t))\right]\leq F_{\lambda_{1}}(u_{*}).$
-From this we find that, see e.g. [8, Inequalities (4.60) and (4.61)],
+From this we find that, see e.g. [9, Inequalities (4.60) and (4.61)],
 $\limsup_{t\to\infty}\left[\max_{x\in\bar{U}}u(x,t)\right]\leq u_{*}.$ (5.59)
@@
 $\mathcal{L}\bar{w}_{2}=\mathcal{L}w_{2}\geq-\mathcal{F}_{2}(t).$
-Consider sufficiently large $t$. By (5.48), we have $m^{\lambda_{2}}\leq\mathcal{F}(t)^{\lambda_{1}}$. Together with (5.54), this yields
+Consider sufficiently large $t$. By (5.48), we have $m^{\lambda_{2}}\leq\mathcal{F}(t)^{\lambda_{1}}$. Together with (5.54) and (5.45), this yields
 $\max_{\Gamma\times[t,t+1]}|\bar{w}_{2}|+\int_{t}^{t+1}\mathcal{F}_{2}(\tau){\rm d}\tau\leq C_{*}\Lambda_{g}(t)+\int_{t}^{t+1}F(\tau)\mathcal{F}(\tau)^{\lambda_{1}}{\rm d}\tau\leq\mu_{5}\widetilde{\Lambda}(t),$
@@
 Step 5. Let $\varepsilon>0$ be arbitrary. By (5.60) and (5.61), we have for sufficiently large $t$ that
-$-\varepsilon\leq\min_{x\in\bar{U}}\bar{u}(x,t)\leq\bar{u}(x,t)\leq\max_{x\in\bar{U}}\bar{u}(x,t)\leq\varepsilon\text{ for all }x\in\bar{U},$
+$\varepsilon\leq\min_{x\in\bar{U}}\bar{u}(x,t)\leq\bar{u}(x,t)\leq\max_{x\in\bar{U}}\bar{u}(x,t)\leq\varepsilon\text{ for all }x\in\bar{U},$
 which implies $|\bar{u}(x,t)|\leq\varepsilon$ for all $x\in\bar{U}$. This proves the limit (5.18). ∎
@@
 Although $P(s)$ is not defined at $s=u_{*}=0$, we have for $\lambda_{1}>0$ the transformation $F_{\lambda_{1}}(s)=\frac{s^{\lambda_{1}+1}}{\lambda_{1}+1}$ for $s>0$ can be extended to the function $\widehat{F}(s)=\frac{s^{\lambda_{1}+1}}{\lambda_{1}+1}$ for all $s\geq 0$.
-Note that $\widehat{F}^{\prime}(s)=s^{\lambda_{1}}$ is bounded near 00.
+Note that $\widehat{F}^{\prime}(s)=s^{\lambda_{1}}$ is bounded near 0.
 In the proof of Theorem 5.9, we replace $F_{\lambda_{1}}$ with $\widehat{F}$ and proceed up to (5.58).
@@
   2. ()
-    For more cases of $u(\overline{U_{\infty}})\subset\bar{J}$ and/or $u_{*}\in\partial J\setminus J$, see [8, Theorem 4.7].
+    For more cases of $u(\overline{U_{\infty}})\subset\bar{J}$ and/or $u_{*}\in\partial J\setminus J$, see [9, Theorem 4.7].
   3. ()
@@
 Data availability. No new data were created or analyzed in this study.
-Funding. No funds were received for conducting this study.
+Funding. A.I. obligatorily acknowledges OGRI’s grant 122022800272-4.
 Conflict of interest. There are no conflicts of interests.